### v2

# FastTrackTr:Towards Fast Multi-Object Tracking with Transformers

Abstract
Transformer-based multi-object tracking (MOT) methods have captured the attention of many researchers in recent years. However, these models often suffer from slow inference speeds due to their structure or other issues. To address this problem, we revisited the Joint Detection and Tracking (JDT) method by looking back at past approaches. By integrating the original JDT approach with some advanced theories, this paper employs an efficient method of information transfer between frames on the DETR, constructing a fast and novel JDT-type MOT framework: FastTrackTr. Thanks to the superiority of this information transfer method, our approach not only reduces the number of queries required during tracking but also avoids the excessive introduction of network structures, ensuring model simplicity. Experimental results indicate that our method has the potential to achieve real-time tracking and exhibits competitive tracking accuracy across multiple datasets.
[Uncaptioned image][Uncaptioned image]
Figure 1: A comparison of the FPS and HOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA 4090. The numerical precision marked with diagonal lines in the right figure is FP32, while that without lines is FP16.
footnotetext: Liao Pan, Yang Feng, Wu Di, Yu Jinwen, Zhao Wenhui and Liu Bo are with Key Laboratory of Information Fusion, School of Automation, Northwestern Polytechnical University,Xi’an,China. (e-mile: liaopan@mail.nwpu.edu.cn,yangfeng@nwpu.edu.cn,wu_di821@mail. nwpu.edu.cn, yujinwen@mail.nwpu.edu.cn, zwh2024202513@mail.nwpu.edu.cn,503790475@qq.com
1 Introduction
Multi-object tracking (MOT) has long been a computer vision task of great interest to researchers and developers, with extensive application prospects across various industries. Currently, the mainstream tracking models are primarily online models, and the paradigm that has garnered widespread attention in recent years within online models remains tracking-by-detection (TBD). This paradigm, originating from SORT [1], has stood the test of time and demonstrates remarkable vitality. Its improved derivative methods [19, 7, 41] continue to be considered state-of-the-art (SOTA).
In recent years, some tracking methods based on transformer [34] have also attracted considerable attention. However, these methods [31, 20, 43, 39, 10, 42, 11] face a significant challenge: suboptimal inference speed. There are various reasons for this, such as the slow performance of the MOTR series, which relies on query-based tracking. The main reason is not the network structure, but issues related to hardware acceleration caused by the variable number of queries (detailed justification can be found in the appendix). Both our experiments and other works [22] have also demonstrated that the inference speed after lightening the network is still not very ideal. Regarding another recent method, MOTIP [11], while it addresses many of MOTR’s issues, its main problem lies in its complex network structure. Even after replacing MOTIP’s Deformable DETR [51] with RT-DETR [48], its inference speed remains unsatisfactory (results shown in Figure 1).
Methods following the tracking-by-detection paradigm often claim to be extremely fast, but in reality, aside from those purely based on kinematics [3], the speed of most MOT systems is not ideal once the time required for the detection model is considered. Purely kinematic models typically perform worse than when they are combined with reid networks. So, is there a way to achieve both fast and accurate tracking? To address this question, we must look back at the JDT (Joint Detection and Tracking) paradigm, which was initially renowned for its speed. Although it has fallen out of favor in recent years due to various reasons, it likely represents a solution to the aforementioned issues. Previously, there was a transformer-based JDT model, known as the TransTrack [31], which shares the same pain points as MOTIP, including having two decoders. Other methods [20, 38, 2], similar to MOTR, track based on queries. CNN-based JDT models [45, 35], on the other hand, typically produce detection results and their corresponding id embeddings for association matching, with their main issue being that the models are outdated, hence they do not keep pace with current standards.
Based on the above discussion, we present FastTrackTr, a Fast multi-object Tracking method grounded in the Transformer architecture and aligned with the JDT paradigm. This tracking method achieves impressive tracking progress through a unique decoder that integrates past information. Additionally, due to the ease of generating masks for this method, the batch size can be set relatively large, making our model considerably more training-friendly compared to some of the latest transformer-based MOT methods. Additionally, due to the ease of generating masks for this method, the batch size can be set relatively large, making our model considerably more training-friendly compared to some of the latest transformer-based MOT methods.
Furthermore, it is noteworthy that models with a single decoder are capable of achieving end-to-end tracking without increasing the number of queries. This point is further discussed in the subsequent sections and the appendix of this document.
In summary, the main innovations of this paper are as follows:
  • 1)
    proposing FastTrackTr, a novel transformer-based MOT method that achieves fast inference speed while maintaining high accuracy.
  • 2)
    introducing a cross-decoder mechanism that implicitly integrates historical trajectory information, eliminating the need for additional queries or decoders.
  • 3)
    Extensive experiments on multiple benchmark datasets, including DanceTrack [32], SportsMOT [6] and MOT17 [21], demonstrate that our model not only achieves competitive accuracy SOTA methods, but also significantly improves inference speed.
2 Related work
2.1 Transformer-based MOT Methods
The earliest Transformer-based MOT methods were TransTrack [31] and TrackFormer [20]. The former is a JDT model, while the latter introduced the concept of track queries, utilizing these special queries to achieve tracking.
Subsequently, MOTR [43] proposed tracklet-aware label assignment (TALA), which optimized label assignment during the training phase to achieve end-to-end tracking. Later improvements to MOTR, such as MOTRv2 [47], MeMOTR [10], MOTRv3 [42], and CO-MOT [39], demonstrated the powerful tracking performance of the MOTR paradigm. Some efforts have been made to address MOTR’s slow inference speed through model lightweighting, but the results have been less than ideal. Recently, alternative Transformer architectures have been proposed. MOTIP’s [11] novel approach of directly predicting tracked object ID numbers is refreshing, while PuTR’s [14] use of Transformers to directly construct a tracking-by-detection paradigm tracker has also been inspirational.
2.2 Tracking-by-Detection
The TBD paradigm can be considered one of the simplest and most practical tracking methods. Since it essentially associates detected targets, this class of methods often yields satisfactory tracking accuracy as long as the detection results are not poor.
The initial SORT algorithm simply utilized the motion state of detected objects for association. Subsequent models like DeepSORT [36] and ByteTrack [46] improved tracking accuracy by enhancing detection precision and employing embedding-based Re-identification (Re-id) networks. Some approaches, like OC-SORT [3] and HybridSORT [41], improved upon SORT using purely kinematics-based methods. Later models such as MotionTrack [24], DeepOC-SORT [19], and SparseTrack [16] demonstrated that simultaneous improvements in both Re-id network utilization and kinematic models yield the most significant performance enhancements.
Currently, there are also algorithms that directly use networks for association. Besides the aforementioned PuTR [14], SUSHI [5] is a tracking method based on graph neural networks.
2.3 Joint Detection and Tracking
Joint detection and tracking aims to achieve detection and tracking simultaneously in a single stage. This approach was common a few years ago but has been seen less frequently in recent years. Here, we briefly review its development history. D&T [8] first proposed a multi-task architecture based on frame-based object detection and cross-frame trajectory regression. Subsequently, Integrated-Detection [52] improved detection performance by combining detection bounding boxes in the current frame with trajectories from the previous frame. Tracktor [30] directly used tracking boxes from the previous frame as candidate regions, then applied bounding box regression to provide tracking boxes for the current step, thus eliminating the box association process. JDE [35] and FairMOT [45] learned object detection and appearance embedding tasks from a shared backbone network. CenterTrack [49] localized targets through tracking-conditioned detection and predicted their offsets from the previous frame. TransTrack used two decoders to handle tracking and detection separately, then matched and associated the results from both decoders. Our FastTrackTr, by introducing a novel decoder, returns to the essence of the JDT method, simplifies the network structure, and achieves fast tracking.
3 Method
3.1 FastTrackTr Architecture
Figure 2: The complete structure of FastTrackTr. For the initial frame, the model functions as a detection system, assigning unique identifiers to each detected object. Subsequently, from the second frame onwards, the historical queries from previous frames and their corresponding masks undergo processing through a single-layer encoder. The output and image feature of this frame are then fed into a historical-cross decoder. Based on the head’s output, detection results and ID embeddings are generated. These are then matched and associated with previous results to produce the final tracking outcome. This architecture achieves the efficient tracking, which can be run in real time on some devices.
The overall structure of FastTrackTr is illustrated in Figure 2. In frame $T_{1}$, a standard decoder is used to initialize the initial historical information. Subsequently, in the following frames, a Historical Encoder is employed to initialize this historical-representative information. In the next frame, a Historical Decoder, which includes an additional cross-attention layer compared to normal decoder layers, processes this historical historical information. The tracking principle is very similar to that of JDT and FairMOT, utilizing an additional ID Embedding head to obtain the object’s appearance embedding, which is then used along with detection results for association and matching. It is worth noting that the backbone network is ResNet [13].
3.2 Learning Detection and ID Embedding
To enable our FastTrackTr to simultaneously output both the target’s position and appearance embedding in a single forward pass, we formulate the problem as follows: Suppose we have a training dataset ${\mathbf{I},\mathbf{B},\mathbf{y}}_{i=1}^{N}$. Here, $\mathbf{I}\in\mathbb{R}^{c\times h\times w}$ represents an image frame, and $\mathbf{B}\in\mathbb{R}^{k\times 4}$ denotes the bounding box annotations for the $k$ targets in this frame. $\mathbf{y}\in\mathbb{Z}^{k}$ represents the partially annotated identity labels, where -1 indicates targets without an identity label. We hope model can output predicted bounding boxes $\hat{\mathbf{B}}\in\mathbb{R}^{\hat{k}\times 4}$ and appearance embeddings $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{k}\times D}$, where $D$ is the dimension of the embedding.
For the output box, this requirement is easily met, as DETR[4] itself is a detection model. To achieve rapid performance, this paper employs RT-DETR [48] as the base model, which inherently possesses state-of-the-art detection capabilities. We have not made significant modifications to its detection component. The loss function for this part remains consistent with RT-DETR, as shown below:
[…]
where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}=$$\{\hat{c},\hat{b}\}$ and $y=\{c,b\},c$ and $b$ represent categories and bounding boxes, respectively. Similar to RT-DETR, we incorporate the IoU score into the objective function of the classification branch, following the approach from VFL [44], to enforce consistency between the classification and localization of positive samples. $\mathcal{L}_{box}$ includes the L1 loss and the generalized IoU loss [25], which are not described here. And, the auxiliary loss from DAB-DETR [15] is also employed in our approach.
The second objective addresses a metric learning problem, specifically the learning of an embedding space where instances of the same identity are close to each other while those of different identities are far apart. Previous methods like JDE [35] and FairMOT [45] used one-hot labels. While this approach is simple and effective, it requires prior knowledge of the number of IDs in the training set to configure the output dimensions of the ID embedding head. This requirement obviously limits the model’s generalization capability and becomes impractical in dense scenarios, as it leads to excessively high embedding dimensions. Moreover, subsequent Transformer-based tracking models predominantly rely on queries for tracking, and the drawbacks of this method have been discussed earlier.
MOTIP [11] employs a predefined ID dictionary and classification methods for object association, whereas PuTR utilizes an association method based on embedding similarity. These two approaches are undoubtedly more generalizable. However, possibly because their decoders primarily process the trajectories of object movements when outputting embeddings, and since the queries used by FastTrackTr to generate embeddings contain more image features and positional information of the current target, these methods fail to enable FastTrackTr to achieve satisfactory convergence.
To address the aforementioned issues, we need to find a method more suitable for image features and with greater generalizability for generating ID embeddings. Therefore, we turn our attention to the training methods of ReID (Re-identification) models. By converting the generation of ID embeddings into a ReID problem, we only need to consider the distances and similarities between embeddings, thereby enhancing the model’s generalization capability.
For the above reasons and for better performance, in this paper, we adopt the more advanced circle loss [33] in this work. Specifically, the id embeddings of the same object between multiple frames is treated as positive pairs denoted as $s_{p}$ and between different objects as negative samples denoted as $s_{n}$, respectively. - $\alpha_{p}$ and $\alpha_{n}$ are the adaptive weighting factors for positive and negative samples, determined by the following equations:
$\alpha_{p}=\left[O_{p}-s_{p}\right]^{+},\quad\alpha_{n}=\left[s_{n}-O_{n}\right]^{+}$ (2)
where $O_{p}=1+m$ and $O_{n}=-m$, with $m$ being a margin hyperparameter. These terms adjust the weight of each sample based on its distance from the ideal similarity value. The margins $\Delta_{p}=1-m$ and $\Delta_{n}=m$ are used to further separate positive and negative pairs in the embedding space.
The logits for positive and negative pairs are then computed as:
$\log_{p}=-\gamma\alpha_{p}(s_{p}-\Delta_{p}),\quad\log_{n}=\gamma\alpha_{n}(s_{n}-\Delta_{n})$ (3)
Finally, the unified Circle Loss function is formulated as follows:
$&\mathcal{L}_{\text{reid}}=\\ &\log\left[1+\sum_{j=1}^{L}\exp\left(\gamma\left(\alpha_{n}^{j}\left(s_{n}^{j}-\Delta_{n}\right)\right)\right)\sum_{i=1}^{K}\exp\left(-\gamma\left(\alpha_{p}^{i}\left(s_{p}^{i}-\Delta_{p}\right)\right)\right)\right]\end{aligned}$
This formulation allows for a dynamic weighting of positive and negative pairs, enhancing the model’s ability to discriminate between similar and dissimilar identities. Circle Loss has proven to be an effective solution for embedding learning in tasks that require fine-grained discrimination, which aligns with the goals of appearance embedding in our tracking framework.
The final total loss is as follows:
$\mathcal{L}_{overall}=\lambda_{reid}*\mathcal{L}_{reid}+\lambda_{det}*\mathcal{L}_{det}$ (5)
where $\lambda_{reid}$ and $\lambda_{det}$ are proportional coefficients used to balance the loss. In this paper, both of them are set to 1. Notably, it should be noted that, when computing the ReID loss, similar to the detection loss, the outputs from each decoder layer are incorporated into the calculation.
3.3 Historical Decoder and Encoder
Historical Decoder: Previous DETR-based MOT models[43, 20, 10] typically feed the final output queries of the previous frame, after some processing, directly into the decoder of the current frame to achieve tracking. This is an explicit tracking method, but it increases the number of queries, thereby increasing computational costs. To more efficiently utilize these historical queries, we replaced the self-attention mechanism in DETR’s decoder with a cross-attention mechanism to process the query features between consecutive frames. This significantly reduces the computational load. The differences between our decoder and the standard decoder are shown in Figure 3.
Specifically, each decoder module in DETR [4] consists of a sequential structure comprising a self-attention layer for the queries, a visual cross-attention layer, and a feed-forward neural network (FFN). Initially, the self-attention layer performs global modeling over the queries to enhance the model’s object detection performance. To further leverage historical information and improve tracking accuracy, we concatenate the final decoder output from frame $t-1$, denoted as $q^{t-1’}_{f}$, which is processed by the trajectory encoder, with the initial queries $q_{d}^{t}$ of frame $t$. This concatenation replaces the key and value in the self-attention layer while keeping the queries unchanged. The implementation can be formally expressed as follows:
$Q_{T}=\operatorname{Linear}(q_{d}^{t}),\quad K_{T}=V_{T}=\operatorname{Linear}(concat(q_{d}^{t},q^{t-1’}_{f}))$ (6)
where $Q_{T}\in\mathbb{R}^{N\times C}$ , $K_{T}\text{ and }V_{T}\in\mathbb{R}^{2N\times C}$. Then, we calculate the query-trajectory attention map $A_{T}\in\mathbb{R}^{2N\times N}$, and aggregate informative depth features weighted by $A_{T}$ to produce the Historical-aware queries $q^{\prime}$, formulated as:
[…]
This mechanism allows each object query to adaptively capture features of the same object from images across consecutive frames, thereby achieving better object understanding. The queries with estimated historical information are then input into the inter-query self-attention layer for feature interaction between objects, and the visual cross-attention layer for collecting visual semantics. We use 6 decoder layers.
Figure 3: Schematic of the flow of our historical-cross decoder and a general decoder. Compared to a general decoder, ours replaces self-attention with historical cross-attention.
Historical Encoder: In FastTrackTr, we have introduced an additional Historical Encoder to enhance the modeling of temporal relationships, and we utilize a masking mechanism during the self-attention process to provide contextual priors for the tracked objects. For the historical feature information from the previous frame or multiple frames, only a subset of the data is utilized. To effectively model these historical data, we employ a masking mechanism and process the data through standard encoder layers. During training, the generation of masks is determined by the ground truth labels, while in inference, masks are based on confidence scores to decide which objects to retain.
The workflow of the historical encoder is illustrated in Figure 4. After generating the mask, we fuse the decoder output $q_{f}^{t}$ with the historical information $q_{f}^{t-1}$ through an attention mechanism, computed as follows:
$Q_{E}=\quad K_{E}=\operatorname{Linear}\left(q^{t-1}_{f}+q_{pos}\right),V_{E}=\operatorname{Linear}(q^{t}_{f})$ (8)
The attention calculation follows the same formula as previously used in the decoder, which is not repeated here. Finally, the fused features are processed by an additional network layer to obtain $q^{t^{\prime}}_{f}$. It is noted that for $t=1$, $q^{t-1}_{f}=q^{t}_{f}$.
The mask generation module in this method includes two modes: (1) In the training phase, the mask is determined by ground truth information, generated by checking if each object matches with the ground truth labels, and random sampling of the masks is performed to flip them, enhancing the model’s robustness; (2) In the inference phase, the mask is determined based on the confidence score of each object, with objects below a certain threshold being filtered out, while high-confidence objects are retained. This masking mechanism ensures that the model focuses on processing significant historical information, thereby improving the encoder’s capability to model the temporal dynamics of target trajectories.
Figure 4: Schematic of the flow of our Historical Encoder. The main role of this Encoder is to further aggregate the temporal features. The generated mask is randomly inverted partially during training. The probability of reversal is related to the number of truth values.
3.4 Matching and Association
Although matching and association and not the primary focus of our paper, it is essential to outline some details of this process. Following the work of JDE, for a given video, after processing each frame, the model outputs detection results along with their corresponding ID embeddings. We then use these values to compute a similarity matrix between the embeddings of the observations and those of the existing tracking chains. The Hungarian algorithm is employed to assign observations to tracking chains. Subsequently, a Kalman filter is used to smooth the trajectories and predict the position of the previous frame’s tracking chain in the current frame. If the assigned observations are spatially distant from the predicted position, the assignment is rejected. Finally, the embeddings of the tracking chains are updated as follows:
$f_{t}=\eta f_{t-1}+(1-\eta)\tilde{f}$ (9)
where $\tilde{f}$ indicates the embedding of the assigned observation, and $f_{t}$ indicates the embedding of the tracklet at frame $t$, $\eta$ is a momentum term for smoothing, and we set $\eta=0.9$.
Of course, considering that there are now many optimized methods for this process, we have also implemented several validated, simple, and effective techniques to enhance this workflow. Specific improvements and their impacts can be found in the appendix.
4 Experiments
Table 1: Overall Latency comparison between FastTrackTr and existing methods on the Dancetrack [32] val. Methods highlighted in gray utilize the YOLOX [12] detector, with its processing time included. Similarly, the latency of our method accounts for its associated time consumption, demonstrating efficient performance with the chosen configuration. $\text{Dec}^{3}$ refers to a 3-layer decoder layer, while R18 refers to ResNet18.
Methods Image Res. Numerical Precision Overall Latency(ms) HOTA
DeepOC-SORT[19] $1440\times 800$ FP16 28.1ms (35.6fps) 58.5
SUSHI[5] $1440\times 800$ FP32/FP16 42.9ms (23.3fps) / 38.6ms (25.9fps) ///
PuTR[14] $1440\times 800$ FP32/FP16 30.5ms (32.8fps) / 28.2ms (35.5fps) 55.7
MOTR [43] $1333\times 800$ FP32/FP16 60.0ms (16.5fps) / 52.9ms (18.9fps) 51.3
MOTIP[11] $1333\times 800$ FP32/FP16 61.7ms (16.2fps) / 45.7ms (21.8fps) 59.1
FastTrackTr $1333\times 800$ FP32/FP16 40.2ms (24.8fps) / 30.9ms (32.4fps) 56.8
MOTIP(RT-DETR) $640\times 640$ FP32/FP16 32.1ms (31.2fps) / 24.5ms (40.8fps) 56.2
FastTrackTr $640\times 640$ FP32/FP16 23.6ms (42.3fps) / 19.6ms (51.0fps) 54.1
FastTrackTr-$\text{Dec}^{3}$ $640\times 640$ FP32/FP16 20.4ms (49.0fps) / 14.8ms (67.6fps) 51.2
FastTrackTr-R18-$\text{Dec}^{3}$ $640\times 640$ FP32/FP16 12.8ms (78.0fps) / 9.86ms (101.4fps) 47.4
4.1 Setting
Datasets: Our proposed method mainly evaluated on DanceTrack [32] and SportsMOT [6], two recently introduced datasets with large-scale training data, which helps prevent overfitting. Both datasets also provide official validation sets, allowing for exploratory experiments. Additionally, we present results on the MOT17 [21] benchmark.
Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA [26, 18]. These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance.
Implementation Details: For inference speed, our FastTrackTr is built on RT-DETR [48], with ResNet50 [13] as the backbone unless specified otherwise. The model is implemented in PyTorch [23], and the training is conducted on 8 NVIDIA RTX 4090 GPUs. The sequence length for training is set to 8 unless otherwise noted.
All ablation experiments use an image size of $640\times 640$. For further testing of the best tracking accuracy of our method, we use an image size of $800\times 1333$ when comparing with SOTA methods. The Trajectory Encoder during training applies random mask transformations, with the probability related to the number of ground-truth objects. Given the number of ground-truth objects $n_{gt}$ and the number of queries $n_{q}$, the probability of a non-masked flip is $\min(0.5,2\cdot n_{gt}/n_{q})$, while for the rest, the probability is $\min(\max(0.1,n_{gt}/n_{q}),0.2)$. For all three datasets, $n_{q}$ is set to 300. Dancetrack and SportMOT are trained for 27 epochs, while MOT17 is trained for 80 epochs. The optimizer is AdwmW [17] with a learning rate of $10^{-4}$ and weight decay of $5.0\times 10^{-4}$. Due to GPU memory limitations, the batch size is set to 4 for smaller images and 1 for larger images. Lastly, for all datasets, the final output dimension of ID embedding is 256.
4.2 Main Result
Comprehensive Comparison of Speed and Accuracy: Table 1 highlights FastTrackTr’s superior latency performance compared to existing methods on the Dancetrack validation set. FastTrackTr consistently delivers faster processing speeds while maintaining competitive HOTA scores, demonstrating efficient real-time capabilities. Additionally, its performance scales effectively with reduced image resolutions, further enhancing its applicability. Advanced variants of FastTrackTr achieve even higher frame rates with only minimal compromises in tracking accuracy. These advantages make FastTrackTr highly effective for practical multi-object tracking applications.
Table 2: Performance comparison between FastTrackTr and existing methods on the Dancetrack [32] test set. The best performance among the methods is marked in bold and please pay more attention to the metrics with *.
Methods HOTA* DetA AssA MOTA IDF1
FairMOT[45] 39.7 66.7 23.8 82.2 40.8
CenterTrack [49] 41.8 78.1 22.6 86.8 35.7
TraDeS [37] 43.3 74.5 25.4 86.2 41.2
TransTrack[31] 45.5 75.9 27.5 88.4 45.2
ByteTrack [46] 47.7 71.0 32.1 89.6 53.9
GTR[50] 48.0 72.5 31.9 84.7 50.3
QDTrack[9] 54.2 80.1 36.8 87.7 50.4
MOTR [43] 54.2 73.5 40.2 79.7 51.5
OC-SORT [3] 55.1 80.3 38.3 92 .0 54.6
PuTR[14] 55.8 /// /// 91.9 58.2
C-BIoU [40] 60.6 81.3 45.4 91.6 61.6
CMTrack [29] 61.8 /// 46.4 92.5 63.3
FastTrackTr(ours) 62.4 78.4 49.8 88.8 64.8
Comparison of State-Of-The-Art: To better demonstrate the superiority of our FastTrackTr, we employed nearly all available data augmentations and used larger image sizes. Tables 2, 3, and 4 respectively showcase the maximum tracking performance of FastTrackTr on DanceTrack, SportsMOT, and MOT17 benchmark. Similar to many previous works [49, 47, 10, 43], we utilized the CrowdHuman [28] dataset for data augmentation on MOT17 and DanceTrack. It is important to note that although our method still does not outperform CNN-based models on the MOT17 benchmark, the underlying reason remains the well-discussed issue: for transformer models, the amount of data in the MOT17 dataset is too limited. Nevertheless, our method has surpassed all previous transformer-based approaches. This indicates that our method performs better even when the dataset images do not allow for sufficient training of the model, suggesting that our approach has potential for practical application in real-world settings, rather than being confined mostly to academic research like previous transformer methods, with few exceptions.
Table 3: Performance comparison with state-of-the-art methods on SportsMOT [6].
Methods HOTA* DetA AssA MOTA IDF1
FairMOT [45] 49.3 70.2 34.7 86.4 53.5
QDTrack [9] 60.4 77.5 47.2 90.1 62.3
ByteTrack [46] 62.1 76.5 50.5 93.4 69.1
OC-SORT [3] 68.1 84.8 54.8 93.4 68.0
MeMOTR [10] 68.8 82.0 57.8 90.2 69.9
FastTrackTr(ours) 70.1 83.9 58.6 94.0 72.0
Table 4: Performance comparison with state-of-the-art methods on MOT17 [21]. The best performance among the transformer-based methods is marked in bold.
Methods HOTA* DetA AssA MOTA IDF1
CNN based:
CenterTrack [49] 52.2 53.8 51.0 67.8 64.7
QDTrack [9] 53.9 55.6 52.7 68.7 66.3
FairMOT [45] 59.3 60.9 58.0 73.7 72.3
ByteTrack [46] 63.1 64.5 62.0 80.3 77.3
OC-SORT [3] 63.2 63.2 63.4 78.0 77.5
C-BIoU [40] 64.1 64.8 63.7 81.1 79.7
MotionTrack [24] 65.1 65.4 65.1 81.1 80.1
Transformer based:
TrackFormer [20] /// /// /// 74.1 68.0
TransTrack [31] 54.1 61.6 47.9 74.5 63.9
TransCenter [38] 54.5 60.1 49.7 73.2 62.2
MeMOT [2] 56.9 /// 55.2 72.5 69.0
MOTR [43] 57.2 58.9 55.8 71.9 68.4
MOTIP [11] 59.2 62.0 56.9 75.5 71.2
PuTR[14] 61.1 /// /// 74.8 74.1
FastTrackTr(our) 62.4 62.8 63.0 76.7 77.2
4.3 Ablation Study
We tested the impact of various components of our model and other factors on performance on the val set of DanceTrack.
The ablation study in Table 5 highlights the progressive enhancements of FastTrackTr on the DanceTrack validation set. Starting with a baseline HOTA of 47.3, adding the Decoder increases HOTA to 51.6, AssA to 37.2, and IDF1 to 51.3. Incorporating both Decoder and Encoder further boosts HOTA to 53.8, AssA to 39.9, and IDF1 to 53.5. Finally, including the Mask module achieves the highest scores with HOTA of 54.1, AssA of 40.0, and IDF1 of 54.7. These results demonstrate that each component significantly contributes to the overall tracking performance of FastTrackTr.
Table 5: Ablation Study of our proposed FastTrackTr on the DanceTrack validation set.
Decoder Encoder Mask HOTA AssA IDF1
47.3 32.5 46.3
$\checkmark$ 51.6 37.2 51.3
$\checkmark$ $\checkmark$ 53.8 39.9 53.5
$\checkmark$ $\checkmark$ $\checkmark$ 54.1 40.0 54.7
In Table 6, we demonstrate the impact of different trajectory cross-attention mechanisms within the trajectory decoder on tracking performance. Here, $q^{t}$ represents the original query, while $q_{f}^{t-1}$ denotes the output of the decoder from the previous frame, with ’concat’ indicating their concatenation. It is important to note that in this table, $q_{f}^{t-1}$ has not passed through the trajectory encoder, hence its performance is not as good as that of the final model.
Table 6: The Impact of Different Historical Cross Attention mechanisms on Tracking Performance
$k$ $v$ HOTA DetA AssA IDF1
$q^{t}$ $q_{f}^{t-1}$ 48.8 70.4 35.8 48.1
$q_{f}^{t-1}$ $q_{f}^{t-1}$ 48.2 70.0 33.0 47.7
$q_{f}^{t-1^{\prime}}$ $q^{t}$ 49.0 71.5 35.9 48.5
concat concat 51.6 72.8 37.2 51.3
Table 7 analyzes the effect of different training frame counts on long-term fusion performance. Utilizing 8 frames yields the highest HOTA, demonstrating superior tracking accuracy. Compared to using fewer frames, 8 frames offer substantial performance improvements while maintaining efficient training times. Beyond 8 frames, metrics show minimal or negligible gains, indicating diminishing returns with additional frames. Thus, selecting 8 training frames provides an optimal balance between enhanced performance and training efficiency for FastTrackTr.
Table 7: Training frames for long-term fusion.
Training frames HOTA DetA AssA IDF1
3 51.7 72.8 36.7 50.9
5 53.5 74.2 38.8 53.5
8 54.1 75.6 40.0 54.7
9 54.0 75.4 39.6 54.8
10 54.1 75.8 39.8 54.6
In Table 8, we demonstrate the impact of using circle loss [33], triplet loss [27], and one-hot encoding loss of the JDE [35] on our method performance during the training of ID embeddings. From the result, circle loss consistently outperforms both triplet loss and one-hot encoding loss across all metrics. Specifically, circle loss achieves the highest tracking performance. In comparison, triplet loss shows the lowest performance, and one-hot encoding loss provides intermediate results. These results demonstrate that circle loss effectively enhances the training of ID embeddings, leading to superior tracking performance in our method.
Table 8: The impact of various ID Embedings losses on performance.
$loss$ HOTA DetA AssA IDF1
triplet loss 52.7 76.8 38.1 52.4
one hot 53.8 75.2 39.4 54.1
circle loss 54.1 75.6 40.0 54.7
5 Conclusion and Limitations
Conclusion: In this paper, we introduce a rapid tracking method, FastTrackTr, which demonstrates superior performance across multiple datasets. By incorporating novel trajectory encoder and decoder designs, we effectively enhance the accuracy and robustness of tracking. Additionally, FastTrackTr achieves significant improvements in computational efficiency, presenting substantial potential for real-time applications. Future work will focus on further optimizing the model structure to adapt to more complex scenarios and exploring deployment possibilities across different hardware platforms.
Limitations: The primary limitation of FastTrackTr is its tracking accuracy, which does not surpass that of MOTIP. This issue likely arises because MOTIP employs a specialized network structure for ID encoding and its training method for long sequences, which ensures better temporal information learning. As the saying goes, ”there are trade-offs”; these aspects also result in poor inference speed and slow training for MOTIP. Conversely, FastTrackTr does not exhibit these drawbacks.
Appendix A More discussion on FastTrackTr
A.1 Description of calculations
In this section, we analyze the computational complexity of FastTrackTr and MOTR. The goal is to determine the threshold at which the computational cost of the MOTR exceeds that of our FastTrackTr when the number of queries increases. We focus on the decoder layer, including the self-attention, cross-attention, feed-forward network (FFN), and normalization layers.
A.1.1 Model Definitions
MOTR (Variable Query Number): The number of queries increases from $N$ to $N_{q}=N+\Delta N$, where $\Delta N$ is the additional number of queries. Both the self-attention and subsequent components are affected by this increase.
FastTrackTr (Modified Self-Attention): The self-attention mechanism is altered such that the key and value tensors have dimensions $2N\times C$, while the query tensor remains $N\times C$. Other components retain their original computational complexity.
A.1.2 Notations
  • $N$: Original number of queries.
  • $\Delta N$: Additional queries in MOTR ($N_{q}=N+\Delta N$).
  • $C$: Dimensionality of each query.
  • $M$: Length of the memory (number of keys/values in cross-attention).
  • $d_{\text{ff}}$: Dimensionality of the FFN’s hidden layer (typically $4C$).
A.1.3 Computational Complexity Analysis
MOTR:
  1. 1.
    Self-Attention:
    $\mathcal{O}_{\text{SA}_{1}}$ $=(N+\Delta N)^{2}\cdot C$ (10)
  2. 2.
    Cross-Attention:
    $\mathcal{O}_{\text{CA}_{1}}$ $=(N+\Delta N)\cdot M\cdot C$ (11)
  3. 3.
    Feed-Forward Network (FFN):
    $\mathcal{O}_{\text{FFN}_{1}}$ $=2(N+\Delta N)\cdot C\cdot d_{\text{ff}}$ (12)
    (Two linear layers with activation in between.)
  4. 4.
    Layer Normalization:
    $\mathcal{O}_{\text{LN}_{1}}$ $=2(N+\Delta N)\cdot C$ (13)
    (Two LayerNorm layers.)
Total Complexity of MOTR:
$\mathcal{O}_{1}$ $=(N+\Delta N)^{2}\cdot C+(N+\Delta N)\cdot M\cdot C$
$\quad+2(N+\Delta N)\cdot C\cdot d_{\text{ff}}+2(N+\Delta N)\cdot C$ (14)
FastTrackTr:
  1. 1.
    Self-Attention:
    $\mathcal{O}_{\text{SA}_{2}}$ $=N\cdot(2N)\cdot C=2N^{2}\cdot C$ (15)
  2. 2.
    Cross-Attention:
    $\mathcal{O}_{\text{CA}_{2}}$ $=N\cdot M\cdot C$ (16)
  3. 3.
    Feed-Forward Network (FFN):
    $\mathcal{O}_{\text{FFN}_{2}}$ $=2N\cdot C\cdot d_{\text{ff}}$ (17)
  4. 4.
    Layer Normalization:
    $\mathcal{O}_{\text{LN}_{2}}$ $=2N\cdot C$ (18)
Total Complexity of FastTrackTr:
$\mathcal{O}_{2}$ $=2N^{2}\cdot C+N\cdot M\cdot C+2N\cdot C\cdot d_{\text{ff}}+2N\cdot C$ (19)
A.1.4 Inequality for Computational Cost Comparison
To find the threshold $\Delta N$ where $\mathcal{O}_{1}>\mathcal{O}_{2}$, we set up the inequality:
$(N+\Delta N)^{2}\cdot C+(N+\Delta N)\cdot M\cdot C$
$\quad+2(N+\Delta N)\cdot C\cdot d_{\text{ff}}+2(N+\Delta N)\cdot C$
$>2N^{2}\cdot C+N\cdot M\cdot C+2N\cdot C\cdot d_{\text{ff}}+2N\cdot C$ (20)
Simplifying the inequality by eliminating common terms and dividing both sides by $C$:
$(N+\Delta N)^{2}-2N^{2}+(N+\Delta N-N)\cdot M$
$\quad+2(N+\Delta N-N)\cdot d_{\text{ff}}+2(N+\Delta N-N)>0$
$\Rightarrow$ $(N+\Delta N)^{2}-2N^{2}+\Delta N\cdot(M+2d_{\text{ff}}+2)>0$ (21)
Expanding the square term:
$N^{2}+2N\Delta N+(\Delta N)^{2}-2N^{2}$
$+\Delta N\cdot(M+2d_{\text{ff}}+2)>0$ (22)
Combining like terms:
$-N^{2}+2N\Delta N+(\Delta N)^{2}$
$+\Delta N\cdot(M+2d_{\text{ff}}+2)>0$ (23)
Simplifying:
$-N^{2}+\Delta N\left(2N+\Delta N+M+2d_{\text{ff}}+2\right)>0$ (24)
A.1.5 Determining the Threshold $\Delta N$
To find the minimum $\Delta N$ satisfying the inequality, we proceed as follows.
$-N^{2}+\Delta N\left(2N+\Delta N+M+2d_{\text{ff}}+2\right)>0$ (25)
A.1.6 Final Solution with Specific Numerical Values
Assuming typical values for a Deformable DETR decoder layer:
  • Number of original queries: $N=300$
  • Memory length: $M=8400$
  • Query dimensionality: $C=256$
  • FFN hidden dimension: $d_{\text{ff}}=4C=1024$
Substituting the numerical values:
$-(300)^{2}+\Delta N\Big{(}2\times 300+\Delta N+8400$
[…]
Computing constants:
$-90,000+\Delta N\left(600+\Delta N+8400+2048+2\right)>0$ (27)
Simplifying the expression inside the parentheses:
$-90,000+\Delta N\left(\Delta N+11,050\right)>0$ (28)
Forming a quadratic inequality:
$(\Delta N)^{2}+11,050\Delta N-90,000>0$ (29)
Solving the quadratic equation:
$(\Delta N)^{2}+11,050\Delta N-90,000=0$ (30)
Using the quadratic formula:
$\Delta N=\frac{-11,050\pm\sqrt{(11,050)^{2}-4\times 1\times(-90,000)}}{2}$ (31)
Calculating the discriminant:
$D$ $=(11,050)^{2}+360,000$
$=122,102,500+360,000=122,462,500$ (32)
Finding the square root:
$\sqrt{D}\approx 11,062.83$ (33)
Determining the positive root:
$\Delta N=\frac{-11,050+11,062.83}{2}\approx\frac{12.83}{2}\approx 6.415$ (34)
The computational cost of MOTR exceeds that of FastTrackTr when the number of additional queries satisfies $\Delta N\geq 7$. This indicates that increasing the number of queries by at least seven will result in higher computational complexity for MOTR compared to FastTrackTr, considering all components of the decoder layer.
In summary, the computational load of FastTrackTr is generally lower than that of MOTR, except in scenarios where there are very few objects to track; in those cases, the theoretical computation may be slightly higher for FastTrackTr. Otherwise, FastTrackTr consistently requires less computational effort. Additionally, in practical applications, MOTR tends to be significantly slower due to memory constraints. For a detailed analysis, please refer to Section. B.
A.2 Possibilities of End-to-end
We originally developed FastTrackTr with the goal of proposing a real-time, end-to-end multi-object tracking model. However, we observed that eliminating one decoder and directly applying MOTIP’s ID loss resulted in significant challenges in model convergence. Upon further analysis, we determined that although MOTIP’s ID loss appears similar to the appearance loss used in the previous JDT models, they fundamentally differ.
Figure 5: The proposed architecture features a hybrid decoder capable of facilitating end-to-end rapid tracking, which currently consists of six layers. There are two distinct types of layers within this decoder: the standard DETR layers, which are used to output detection results, and the ID decoder layers, which are dedicated to producing tracking outcomes. These layers are arranged alternately, with each type comprising three layers. This design aims to effectively merge detection and tracking functionalities within a single streamlined framework.
For models like JDE [35], the focus is on the similarity of object appearances, which does not incorporate specific object trajectories. In contrast, the ID decoder in MOTIP primarily handles trajectory information, with image data serving only as a supplementary element. The function and training method of this decoder might be similar to the entire PuTR model. Although our FastTrackTr attempts to integrate some temporal information by passing queries from one frame to the next, these queries predominantly contain image features. Under such conditions, supervising motion trajectories directly using MOTIP’s method becomes less effective, leading to poor or nearly impossible model convergence.
Following this analysis, we attempted to incorporate more historical information into the decoder. Our preliminary design introduces an identity-specific decoder layer into the traditional DETR architecture, as illustrated in Figure.5. These layers enhance alignment between current frame detections and corresponding identity features through a two-step attention mechanism, and link current detections with historical trajectories by emphasizing temporal consistency. The ID decoder layers are interlaced with standard DETR layers, initially comprising six layers in total.
The trajectory attention within the ID decoder layers can be formally described as follows:
[…]
where $emb_{t}$ and the input id decoder in MOTIP are the same. Its main components are the output of the previous decoder layer and the id encoding. $emb_{h}$ is the historical encoding. Finally, the $q^{\prime}_{ID}\in\mathbb{R}^{2\times N\times C}$ will be reduced in dimension and fused through a conv layer and an FFN layer in the feedforward network.
The id loss used after the feedforward neural network and the id head is the id loss of MOTIP, which is shown below:
$&\mathcal{L}_{id}=\frac{-\sum_{t=2}^{T+1}\sum_{m=1}^{M_{t}}\sum_{k=1}^{K+1}y_{m}^{k}\log\left(p_{m}^{k}\right)}{\sum_{t=2}^{T+1}M_{t}},\\ &y_{m}^{k}=\begin{cases}1&\mathrm{gt}_{id}\text{ of }m^{th}\text{ object is }k,\\ 0&\text{ else },\end{cases}\end{aligned}$ (36)
where $T+1$ denotes the number of frames in a training clip, while there are $M_{t}$ ground-truth objects in the $t$-th frame. $y_{m}^{k}$ is an indicator function according to the identity ground truth of each object, as shown in Eq.(37) . In practice, we simultaneously train object detector and ID predictor in an end-to-end strategy. Therefore, we leverage an overall loss function $\mathcal{L}$ to supervise these two parts:
$\mathcal{L}=\lambda_{\text{det}}\mathcal{L}_{\text{det}}+\lambda_{\text{id }}\mathcal{L}_{\text{id }}$ (37)
Due to time constraints and the incomplete state of the model, we have conducted only one epoch of training on the DanceTrack dataset to observe whether the loss converges. After this single epoch, when employing the MOTIP process during inference, the HOTA score was approximately 16.7. This result preliminarily confirms that the model is broadly effective, with further fine-tuning required to optimize its performance.
A.3 Improvements of the JDE association module
We believe that one of the reasons why JDT-type models have gradually been abandoned in recent years is that the commonly used JDE [35] association module has become somewhat outdated. This has led to JDT models underperforming compared to some of the more recent models. To better demonstrate the superiority of our approach, we made some minor adjustments to the JDE association method to improve its tracking performance.
Our improvements mainly focus on two aspects:
  • 1)
    In the JDE association module, we associate only high-confidence detection boxes, ignoring those with lower confidence scores. By incorporating ByteTrack’s [46] bi-stage matching mechanism, we temporarily retain these low-confidence detection boxes as candidate targets. After matching the high-confidence targets, we perform a second round of matching using the high-confidence targets, which enhances tracking performance to some extent.
  • 2)
    We also made some adjustments to the association strategy. In JDE, the first association is typically based on ID features, followed by a second association using historical trajectories. In our approach, similar to many other trackers, we include historical trajectories in the first round of association. Specifically, the cost matrix $C$ is the weighted sum of the appearance cost $A_{a}$ and the motion cost $A_{m}$, as follows:
Figure 6: The changes in GPU memory consumption after the decoder for the MOTR, Deformable DETR, and our improved model in MOT17-03.
Figure 7: A simple MOTR and our improved tracking process diagram. The number of queries for MOTR will change, but after our improvement, it remains unchanged.
$C=\lambda A_{a}+(1-\lambda)A_{m}$ (38)
where the weight factor $\lambda$ is set to 0.99 , as in [19, 41] .
Table 9: Ablation Study of association module on the DanceTrack validation set.
Byte Cost HOTA DetA AssA IDF1
51.5 72.4 37.5 51.4
$\checkmark$ 53.4 74.3 39.7 53.1
$\checkmark$ 52.4 73.1 38.3 52.6
$\checkmark$ $\checkmark$ 54.1 75.6 40.0 54.7
Table 10: Impact of different of positional encodings of the historical encode on the DanceTrack validation set.
Mechanism HOTA DetA AssA IDF1
w///o 53.8 74.9 39.5 53.8
MLP 53.5 74.2 38.9 53.5
decoder 54.1 75.6 40.0 54.7
A.4 More Implementation Details
By default, on DanceTrack [32], we train FastTrackTr for 27 epochs on the training set, dropping the learning rate by a factor of 10 at the 14th and 18th epochs. On SportsMOT [6], we train our model for 27 epochs on the training set, with the learning rate dropped at the 14th and 18th epochs. On MOT17 [21], we train the model on this combined training set for 80 epochs, with the learning rate dropped at the 45th and 55th epochs.
For the additional use of the CrowdHuman dataset on DanceTrack and MOT17, our usage method follows the same approach as MOTIP. Interested readers can refer to the original work for further details.
Lastly, regarding the positional encoding of the historical encoder, the Table 10 shows the impact of different positional encodings on performance. In this table, ”w/o” denotes the absence of positional encoding, ”MLP” indicates that the positional encoding is generated from an MLP, and ”decoder” refers to using the positional encoding from the final layer of the decoder.
Appendix B How to Improve MOTR inference Speed
[…]
This led us to speculate whether the issue might be due to the dynamic nature of the query system, which could result in variable memory usage. As shown in Figure 6, we compare the memory consumption of the decoder when using fixed queries versus dynamic queries. After further investigation, we found that this is indeed a plausible cause. In PyTorch, memory is managed through cudaMalloc and cudaFree, and this dynamic memory management approach could lead to frequent memory allocation calls in models like MOTR, where memory demand fluctuates. This could slow down the model and prevent full GPU utilization.
We experimented with PyTorch’s memory management code, and observed that when the tensor size remains relatively constant, there are significantly fewer memory allocation and release calls. The specific code we tested can be easily found in GitHub, when you search for ’torch mem’(we would like to note that the owner of that link is not associated with the authors of this paper.
So how can we address this issue? Modifying PyTorch’s underlying logic directly is quite complex and would likely require intervention from PyTorch’s developers (we are preparing to submit an issue on GitHub or send an email to the maintainers). An alternative approach could be to fix the number of tracking queries in MOTR and use masks to differentiate between empty tracking queries and those that are actively used. This may resolve the issue. As shown in Figure 7, we trained an updated version of MOTR for one epoch on the DanceTrack dataset and tested the speed and GPU utilization. Due to limitations in funding, equipment, and time, we did not fully train the model to evaluate its performance. However, preliminary results show promise: after one epoch, the HOTA score on the validation set reached 18.5 and fps has increased from 16.5 to 25 with fp32 precision. Note that we have not reduced any network structure at this time. Interested readers are welcome to experiment with this approach themselves. It is relatively straightforward to implement, though we advise against using GPUs with less than 24GB of memory, as this may lead to out-of-memory errors.
Of course, this is just one possible explanation. Other factors may also contribute to the underutilization of GPU resources in MOTR-like models. If any readers have alternative insights, we welcome further discussion and collaboration.
Appendix C Visualization of FastTrackTr Results
In this section, we will provide visualization results of our model on different datasets.
Figure 8: Our model’s visualization results on Dancetrack, where the same color represents the same object.
Figure 9: Our model’s visualization results on MOT17, where the same color represents the same object.
Figure 10: Our model’s visualization results on SportsMOT, where the same color represents the same object.

### v3

@@
-FastTrackTr:Towards Fast Multi-Object Tracking with Transformers
+FastTrackTr:Towards Real-time Multi-Object Tracking with Transformers
 Abstract
@@
 Experimental results indicate that our method has the potential to achieve real-time tracking and exhibits competitive tracking accuracy across multiple datasets.
-[Uncaptioned image][Uncaptioned image]
-Figure 1: A comparison of the FPS and HOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA 4090.
-The numerical precision marked with diagonal lines in the right figure is FP32, while that without lines is FP16.
-footnotetext: Liao Pan, Yang Feng, Wu Di, Yu Jinwen, Zhao Wenhui and Liu Bo are with Key Laboratory of Information Fusion, School of Automation, Northwestern Polytechnical University,Xi’an,China.
-(e-mile: liaopan@mail.nwpu.edu.cn,yangfeng@nwpu.edu.cn,wu_di821@mail.
-nwpu.edu.cn, yujinwen@mail.nwpu.edu.cn, zwh2024202513@mail.nwpu.edu.cn,503790475@qq.com
+Figure 1: A comparison of the FPS, HOTA and MOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA 4090.
+For the left image, the horizontal axis is FPS, the vertical axis is HOTA, and the radius of the shape is MOTA.
 1 Introduction
@@
 This paradigm, originating from SORT [2], has stood the test of time and demonstrates remarkable vitality.
-Its improved derivative methods [19, 7, 41] continue to be considered state-of-the-art (SOTA).
+Its improved derivative methods [21, 8, 43, 1] continue to be considered state-of-the-art (SOTA).
 In recent years, some tracking methods based on transformer [36] have also attracted considerable attention.
@@
 Even after replacing MOTIP’s Deformable DETR [53] with RT-DETR [50], its inference speed remains unsatisfactory (results shown in Figure 1).
+And, during its implementation, some of the data is dynamic, which makes it relatively difficult to accelerate using TensorRT.
 Methods following the tracking-by-detection paradigm often claim to be extremely fast, but in reality, aside from those purely based on kinematics [4], the speed of most MOT systems is not ideal once the time required for the detection model is considered.
@@
 Additionally, due to the ease of generating masks for this method, the batch size can be set relatively large, making our model considerably more training-friendly compared to some of the latest transformer-based MOT methods.
-Furthermore, it is noteworthy that models with a single decoder are capable of achieving end-to-end tracking without increasing the number of queries.
-This point is further discussed in the subsequent sections and the appendix of this document.
+Furthermore, the deterministic data architecture inherent in FastTrackTr enables seamless TensorRT acceleration, achieving genuine real-time performance.
+This computational efficiency remains largely unattainable for contemporary transformer-based MOT methods due to their inherent dynamic computation patterns.
 In summary, the main innovations of this paper are as follows:
   • 1)
-    proposing FastTrackTr, a novel transformer-based MOT method that achieves fast inference speed while maintaining high accuracy.
+    proposing FastTrackTr, a novel transformer-based MOT method that achieves real-time inference speed while maintaining high accuracy.
   • 2)
@@
 Later models such as MotionTrack [26], DeepOC-SORT [21], and SparseTrack [18] demonstrated that simultaneous improvements in both Re-id network utilization and kinematic models yield the most significant performance enhancements.
-Currently, there are also algorithms that directly use networks for association. Besides the aforementioned PuTR [14], SUSHI [5] is a tracking method based on graph neural networks.
+Currently, there are also algorithms that directly use networks for association.
+Besides the aforementioned PuTR [16], SUSHI [6] is a tracking method based on graph neural networks.
 2.3 Joint Detection and Tracking
@@
 Subsequently, in the following frames, a Historical Encoder is employed to initialize this historical-representative information.
-In the next frame, a Historical Decoder, which includes an additional cross-attention layer compared to normal decoder layers, processes this historical historical information.
+In the next frame, a Historical Decoder, which includes a historical cross-attention layer compared to normal decoder layers, processes this historical information.
 The tracking principle is very similar to that of JDT and FairMOT, utilizing an additional ID Embedding head to obtain the object’s appearance embedding, which is then used along with detection results for association and matching.
@@
 The second objective addresses a metric learning problem, specifically the learning of an embedding space where instances of the same identity are close to each other while those of different identities are far apart.
-Previous methods like JDE [35] and FairMOT [45] used one-hot labels.
-While this approach is simple and effective, it requires prior knowledge of the number of IDs in the training set to configure the output dimensions of the ID embedding head.
-This requirement obviously limits the model’s generalization capability and becomes impractical in dense scenarios, as it leads to excessively high embedding dimensions.
-Moreover, subsequent Transformer-based tracking models predominantly rely on queries for tracking, and the drawbacks of this method have been discussed earlier.
-MOTIP [11] employs a predefined ID dictionary and classification methods for object association, whereas PuTR utilizes an association method based on embedding similarity.
-These two approaches are undoubtedly more generalizable.
-However, possibly because their decoders primarily process the trajectories of object movements when outputting embeddings, and since the queries used by FastTrackTr to generate embeddings contain more image features and positional information of the current target, these methods fail to enable FastTrackTr to achieve satisfactory convergence.
-To address the aforementioned issues, we need to find a method more suitable for image features and with greater generalizability for generating ID embeddings.
-Therefore, we turn our attention to the training methods of ReID (Re-identification) models.
-By converting the generation of ID embeddings into a ReID problem, we only need to consider the distances and similarities between embeddings, thereby enhancing the model’s generalization capability.
+Previous JDT methods (e.g., JDE [37], FairMOT [47]) rely on one-hot labels to construct ID embeddings, necessitating predefined ID quantities to configure embedding dimensions.
+This design introduces two critical limitations: (1) explosive dimensionality growth in dense scenes and (2) severely restricted generalization due to dataset-specific ID dependencies.
+Subsequent Transformer-based trackers inherit these constraints through their query mechanisms.
+While recent approaches like MOTIP [12] (using ID dictionaries) and PuTR (leveraging embedding similarity) improve generalizability, their trajectory decoders overemphasize motion patterns, conflicting with FastTrackTr’s queries that encode richer spatial features and positional priors.
+This mismatch hinders effective convergence.
+To resolve these issues, we reformulate ID embedding learning as a re-identification (ReID) problem, decoupling embedding dimensions from ID quantities and prioritizing feature distance/similarity optimization.
+This paradigm shift enhances cross-scene adaptability and better aligns with image-driven tracking frameworks.
 For the above reasons and for better performance, in this paper, we adopt the more advanced circle loss [35] in this work.
-Specifically, the id embeddings of the same object between multiple frames is treated as positive pairs denoted as $s_{p}$ and between different objects as negative samples denoted as $s_{n}$, respectively.
+Specifically, by utilizing the dot product similarity, the id embeddings of the same object across multiple frames are regarded as positive pairs indicated as $s_{p}$, while those between distinct objects are treated as negative samples represented by $s_{n}$.
 - $\alpha_{p}$ and $\alpha_{n}$ are the adaptive weighting factors for positive and negative samples, determined by the following equations:
@@
 Initially, the self-attention layer performs global modeling over the queries to enhance the model’s object detection performance.
-To further leverage historical information and improve tracking accuracy, we concatenate the final decoder output from frame $t-1$, denoted as $q^{t-1’}_{f}$, which is processed by the trajectory encoder, with the initial queries $q_{d}^{t}$ of frame $t$.
+To further leverage historical information and improve tracking accuracy, we concatenate the Historical Encoder output from frame $t-1$, denoted as $q^{t-1’}_{f}$, which is processed by the trajectory encoder, with the initial queries $q_{d}^{t}$ of frame $t$.
 This concatenation replaces the key and value in the self-attention layer while keeping the queries unchanged.
@@
 The queries with estimated historical information are then input into the inter-query self-attention layer for feature interaction between objects, and the visual cross-attention layer for collecting visual semantics.
-We use 6 decoder layers.
 Figure 3: Schematic of the flow of our historical-cross decoder and a general decoder. Compared to a general decoder, ours replaces self-attention with historical cross-attention.
@@
 The workflow of the historical encoder is illustrated in Figure 4.
-After generating the mask, we fuse the decoder output $q_{f}^{t}$ with the historical information $q_{f}^{t-1}$ through an attention mechanism, computed as follows:
-$Q_{E}=\quad K_{E}=\operatorname{Linear}\left(q^{t-1}_{f}+q_{pos}\right),V_{E}=\operatorname{Linear}(q^{t}_{f})$ (8)
-The attention calculation follows the same formula as previously used in the decoder, which is not repeated here.
-Finally, the fused features are processed by an additional network layer to obtain $q^{t^{\prime}}_{f}$.
-It is noted that for $t=1$, $q^{t-1}_{f}=q^{t}_{f}$.
-The mask generation module in this method includes two modes: (1) In the training phase, the mask is determined by ground truth information, generated by checking if each object matches with the ground truth labels, and random sampling of the masks is performed to flip them, enhancing the model’s robustness; (2) In the inference phase, the mask is determined based on the confidence score of each object, with objects below a certain threshold being filtered out, while high-confidence objects are retained.
-This masking mechanism ensures that the model focuses on processing significant historical information, thereby improving the encoder’s capability to model the temporal dynamics of target trajectories.
+After generating the mask, we fuse the decoder output $q_{f}^{t}$ with the historical information $q_{f}^{t-1^{\prime}}$ through an attention mechanism, computed as follows:
+$Q_{E}=K_{E}=\operatorname{Linear}(q^{t}_{f}+q_{pos}),V_{E}=\operatorname{Linear}(q^{t-1^{\prime}}_{f})$ (8)
+where $q_{pos}$ represents a set of learnable positional encodings and the attention weights are computed via scaled dot-product as previously used in the decoder, followed by masking to suppress less relevant historical elements.
+Finally, the fused features are then processed by a feed-forward network to obtain $q^{t^{\prime}}_{f}$.
+For $t=1$, we initialize $q^{t-1^{\prime}}_{f}$ as $q^{t}_{f}$ due to the absence of prior history.
+To prioritize significant historical information, we design a dynamic masking strategy with two operational modes: (1) In the training phase, the mask is determined by ground truth information, generated by checking if each object matches with the ground truth labels, and random sampling of the masks is performed to flip them, enhancing the model’s robustness; (2) In the inference phase, the mask is determined based on the confidence score of each object, with objects below a certain threshold being filtered out, while high-confidence objects are retained.
+The mask adopts [36]’s softmax suppression via large negative biases to irrelevant entries.
+It strengthens the encoder’s capability to model temporal dependencies while suppressing error propagation from noisy historical inputs.
 Figure 4: Schematic of the flow of our Historical Encoder.
@@
 Table 1: Overall Latency comparison between FastTrackTr and existing methods on the Dancetrack [34] val.
-Methods highlighted in gray utilize the YOLOX [12] detector, with its processing time included.
-Similarly, the latency of our method accounts for its associated time consumption, demonstrating efficient performance with the chosen configuration.
-$\text{Dec}^{3}$ refers to a 3-layer decoder layer, while R18 refers to ResNet18.
-Methods Image Res. Numerical Precision Overall Latency(ms) HOTA
-DeepOC-SORT[19] $1440\times 800$ FP16 28.1ms (35.6fps) 58.5
-SUSHI[5] $1440\times 800$ FP32/FP16 42.9ms (23.3fps) / 38.6ms (25.9fps) ///
-PuTR[14] $1440\times 800$ FP32/FP16 30.5ms (32.8fps) / 28.2ms (35.5fps) 55.7
-MOTR [43] $1333\times 800$ FP32/FP16 60.0ms (16.5fps) / 52.9ms (18.9fps) 51.3
-MOTIP[11] $1333\times 800$ FP32/FP16 61.7ms (16.2fps) / 45.7ms (21.8fps) 59.1
-FastTrackTr $1333\times 800$ FP32/FP16 40.2ms (24.8fps) / 30.9ms (32.4fps) 56.8
-MOTIP(RT-DETR) $640\times 640$ FP32/FP16 32.1ms (31.2fps) / 24.5ms (40.8fps) 56.2
-FastTrackTr $640\times 640$ FP32/FP16 23.6ms (42.3fps) / 19.6ms (51.0fps) 54.1
-FastTrackTr-$\text{Dec}^{3}$ $640\times 640$ FP32/FP16 20.4ms (49.0fps) / 14.8ms (67.6fps) 51.2
-FastTrackTr-R18-$\text{Dec}^{3}$ $640\times 640$ FP32/FP16 12.8ms (78.0fps) / 9.86ms (101.4fps) 47.4
+* denotes that these methods employ the same YOLOX [13] detector, and the latency of this detector is accounted for in this table.
+Similarly, the latency of our method accounts for its associated time consumption, and these are highlighted in color.
+$\text{Dec}^{3}$ refers to a 3-layer decoder, while R18 refers to ResNet18.
+Unless otherwise specified, all Transformer-based models have ResNet50 for Backbone and 6 layers for decoder.
+Methods Image Res. Numerical Precision Overall Latency(ms) HOTA Overall Latency with TensorRT FP16(ms)
+Deep OC-SORT*[21] 1333×800 FP16 36.6ms (27.8fps) 58.5 13.5ms(74.1fps)
+SUSHI*[6] 1333×800 FP32/FP16 42.7ms (23.4fps) / 38.0ms (26.3fps) 58.7 -
+PuTR*[16] 1333×800 FP32/FP16 29.9ms (33.4fps) / 27.7ms (36.1fps) 54.5 -
+MOTR[45] 1333×800 FP32/FP16 60.0ms (16.5fps) / 52.9ms (18.9fps) 51.3 -
+MOTIP[12] 1333×800 FP32/FP16 61.7ms (16.2fps) / 45.7ms (21.8fps) 59.1 -
+FastTrackTr 1333×800 FP32/FP16 40.2ms (24.8fps) / 30.9ms (32.4fps) 56.8 11.6ms(86.6fps)
+MOTIP(RT-DETR) 640×640 FP32/FP16 32.1ms (31.2fps) / 24.5ms (40.8fps) 56.2 -
+FastTrackTr 640×640 FP32/FP16 23.6ms (42.3fps) / 19.6ms (51.0fps) 54.1 6.0ms(166.4fps)
+FastTrackTr-Dec3 640×640 FP32/FP16 20.4ms (49.0fps) / 14.8ms (67.6fps) 51.2 4.9ms(203.7fps)
+FastTrackTr-R18-Dec3 640×640 FP32/FP16 12.8ms (78.0fps) / 9.86ms (101.4fps) 47.4 4.1ms(243.4fps)
 4.1 Setting
@@
 For further testing of the best tracking accuracy of our method, we use an image size of $800\times 1333$ when comparing with SOTA methods.
-The Trajectory Encoder during training applies random mask transformations, with the probability related to the number of ground-truth objects.
+The Historical Encoder during training applies random mask transformations, with the probability related to the number of ground-truth objects.
 Given the number of ground-truth objects $n_{gt}$ and the number of queries $n_{q}$, the probability of a non-masked flip is $\min(0.5,2\cdot n_{gt}/n_{q})$, while for the rest, the probability is $\min(\max(0.1,n_{gt}/n_{q}),0.2)$.
@@
 Lastly, for all datasets, the final output dimension of ID embedding is 256.
+All models are implemented in PyTorch by default unless explicitly stated otherwise.
+The complete codebase remains exclusively Python throughout our experiments.
 4.2 Main Result
-Comprehensive Comparison of Speed and Accuracy: Table 1 highlights FastTrackTr’s superior latency performance compared to existing methods on the Dancetrack validation set.
-FastTrackTr consistently delivers faster processing speeds while maintaining competitive HOTA scores, demonstrating efficient real-time capabilities.
-Additionally, its performance scales effectively with reduced image resolutions, further enhancing its applicability.
-Advanced variants of FastTrackTr achieve even higher frame rates with only minimal compromises in tracking accuracy.
-These advantages make FastTrackTr highly effective for practical multi-object tracking applications.
+Comprehensive Comparison of Speed and Accuracy: In RT-DETR, a standardized benchmark has been established for real-time detectors.
+Within this benchmark, all models undergo acceleration via TensorRT, with computational latency measurements exclusively encompassing inference and post-processing durations.
+This paper adopts the same configuration while substituting the detection dataset COCO2017 [15] with the MOT dataset Dancetrack .
+However, this substitution reveals a critical implementation challenge: most transformer-based MOT methods [33, 22, 45, 41, 11, 44, 12, 16] developed in PyTorch incorporate dynamic computational graphs during inference, presumably to reduce computational overhead or simplify implementation.
+These dynamic elements create significant compatibility issues (specifically non-fixed tensor shapes and conditional branching) when converting models to TensorRT - challenges that remain unresolved in our current implementation.
+To ensure fair comparison, Table 1 comprehensively presents both speed and accuracy metrics across tracking methods using various numerical precisions in PyTorch.
+Notably, for our proposed models and the TBD method, only the network benefits from TensorRT acceleration, while the association module retains its original Python implementation.
 Table 2: Performance comparison between FastTrackTr and existing methods on the Dancetrack [34] test set.
@@
 OC-SORT[4] 55.1 80.3 38.3 92.0 54.6
+MoveSORT[1] 56.1 81.6 38.7 91.8 56.0
 PuTR[16] 55.8 /// /// 91.9 58.2
@@
 Comparison of State-Of-The-Art: To better demonstrate the superiority of our FastTrackTr, we employed nearly all available data augmentations and used larger image sizes.
-Tables 2, 3, and 4 respectively showcase the maximum tracking performance of FastTrackTr on DanceTrack, SportsMOT, and MOT17 benchmark.
+Tables 2, 4, and 3 respectively showcase the maximum tracking performance of FastTrackTr on DanceTrack, SportsMOT, and MOT17 benchmark.
 Similar to many previous works [51, 49, 11, 45], we utilized the CrowdHuman [30] dataset for data augmentation on MOT17 and DanceTrack.
@@
 This indicates that our method performs better even when the dataset images do not allow for sufficient training of the model, suggesting that our approach has potential for practical application in real-world settings, rather than being confined mostly to academic research like previous transformer methods, with few exceptions.
-Table 3: Performance comparison with state-of-the-art methods on SportsMOT [6].
-Methods HOTA* DetA AssA MOTA IDF1
-FairMOT [45] 49.3 70.2 34.7 86.4 53.5
-QDTrack [9] 60.4 77.5 47.2 90.1 62.3
-ByteTrack [46] 62.1 76.5 50.5 93.4 69.1
-OC-SORT [3] 68.1 84.8 54.8 93.4 68.0
-MeMOTR [10] 68.8 82.0 57.8 90.2 69.9
-FastTrackTr(ours) 70.1 83.9 58.6 94.0 72.0
 Table 3: Performance comparison with state-of-the-art methods on MOT17 [23]. The best performance among the transformer-based methods is marked in bold.
@@
 PuTR[16] 61.1 /// /// 74.8 74.1
-FastTrackTr(our) 62.4 62.8 63.0 76.7 77.2
+FastTrackTr(ours) 62.4 62.8 63.0 76.7 77.2
+Table 4: Performance comparison with state-of-the-art methods on SportsMOT [7].
+Methods HOTA* DetA AssA MOTA IDF1
+FairMOT [47] 49.3 70.2 34.7 86.4 53.5
+QDTrack [10] 60.4 77.5 47.2 90.1 62.3
+ByteTrack [48] 62.1 76.5 50.5 93.4 69.1
+OC-SORT [4] 68.1 84.8 54.8 93.4 68.0
+MeMOTR [11] 68.8 82.0 57.8 90.2 69.9
+FastTrackTr(ours) 70.1 83.9 58.6 94.0 72.0
 4.3 Ablation Study
@@
 Here, $q^{t}$ represents the original query, while $q_{f}^{t-1}$ denotes the output of the decoder from the previous frame, with ’concat’ indicating their concatenation.
-It is important to note that in this table, $q_{f}^{t-1}$ has not passed through the trajectory encoder, hence its performance is not as good as that of the final model.
+It is important to note that in this table, $q_{f}^{t-1}$ has not passed through the historical encoder, hence its performance is not as good as that of the final model.
 Table 6: The Impact of Different Historical Cross Attention mechanisms on Tracking Performance
@@
 $q_{f}^{t-1}$ $q_{f}^{t-1}$ 48.2 70.0 33.0 47.7
-$q_{f}^{t-1^{\prime}}$ $q^{t}$ 49.0 71.5 35.9 48.5
+$q_{f}^{t-1}$ $q^{t}$ 49.0 71.5 35.9 48.5
 concat concat 51.6 72.8 37.2 51.3
-Table 7 analyzes the effect of different training frame counts on long-term fusion performance.
+Table 7 analyzes the effect of different length of video clips during training.
 Utilizing 8 frames yields the highest HOTA, demonstrating superior tracking accuracy.
@@
 Thus, selecting 8 training frames provides an optimal balance between enhanced performance and training efficiency for FastTrackTr.
-Table 7: Training frames for long-term fusion.
+Table 7: The length of video clips during training.
 Training frames HOTA DetA AssA IDF1
@@
 After further investigation, we found that this is indeed a plausible cause.
-In PyTorch, memory is managed through cudaMalloc and cudaFree, and this dynamic memory management approach could lead to frequent memory allocation calls in models like MOTR, where memory demand fluctuates.
-This could slow down the model and prevent full GPU utilization.
+In cuda, memory is managed through cudaMalloc and cudaFree.
+These operations incur significant computational overhead.
+To mitigate this, PyTorch attempts to reuse memory blocks previously allocated via cudaMalloc.
+When a suitable cached block exists in PyTorch’s allocator, it directly serves the request without invoking cudaMalloc, thereby minimizing latency during initial allocation phases.
+However, MOTR’s dynamic data patterns require intermediate tensors of varying sizes throughout runtime.
+This variability prevents the allocator from leveraging cached blocks of appropriate dimensions, forcing frequent calls to cudaFree to release outdated allocations and subsequent cudaMalloc requests for new memory spaces.
+Consequently, the allocator repeatedly rebuilds its caching infrastructure through bulk CUDA API operations – an expensive process that introduces substantial latency spikes.
+This thrashing behavior substantially degrades computational efficiency and hinders full GPU utilization
 We experimented with PyTorch’s memory management code, and observed that when the tensor size remains relatively constant, there are significantly fewer memory allocation and release calls.

### v4

@@
-FastTrackTr:Towards Real-time Multi-Object Tracking with Transformers
+FastTrackTr: Real-time Multi-Object Tracking with Transformers for Real World
 Abstract
-Transformer-based multi-object tracking (MOT) methods have captured the attention of many researchers in recent years.
-However, these models often suffer from slow inference speeds due to their structure or other issues.
-To address this problem, we revisited the Joint Detection and Tracking (JDT) method by looking back at past approaches.
-By integrating the original JDT approach with some advanced theories, this paper employs an efficient method of information transfer between frames on the DETR, constructing a fast and novel JDT-type MOT framework: FastTrackTr.
-Thanks to the superiority of this information transfer method, our approach not only reduces the number of queries required during tracking but also avoids the excessive introduction of network structures, ensuring model simplicity.
-Experimental results indicate that our method has the potential to achieve real-time tracking and exhibits competitive tracking accuracy across multiple datasets.
-Figure 1: A comparison of the FPS, HOTA and MOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA 4090.
-For the left image, the horizontal axis is FPS, the vertical axis is HOTA, and the radius of the shape is MOTA.
+Transformer-based multi-object tracking (MOT) methods have attracted significant attention from researchers.
+However, these Transformer-based models often suffer from suboptimal inference speeds due to their architectural complexities or other inherent issues, rendering them difficult to deploy in practical industrial applications.
+To address this challenge, we revisited the classic joint detection and tracking (JDT) paradigm and analyzed existing models.
+Drawing inspiration from DETR’s object queries, which naturally encode object appearance features, we constructed a fast and novel JDT-type MOT framework named FastTrackTr by implementing an efficient inter-frame information transfer mechanism.
+Benefiting from the superiority of this mechanism, our approach not only reduces the number of queries required for tracking but also avoids introducing excessive network structures, ensuring model simplicity.
+Experimental results demonstrate that our method not only achieves competitive tracking accuracy across multiple datasets but also enables real-time tracking in practical application environments.
+{IEEEkeywords}
+Multi-object tracking, Transformers, Real-time
 1 Introduction
-Multi-object tracking (MOT) has long been a computer vision task of great interest to researchers and developers, with extensive application prospects across various industries.
-Currently, the mainstream tracking models are primarily online models, and the paradigm that has garnered widespread attention in recent years within online models remains tracking-by-detection (TBD).
-This paradigm, originating from SORT [2], has stood the test of time and demonstrates remarkable vitality.
-Its improved derivative methods [21, 8, 43, 1] continue to be considered state-of-the-art (SOTA).
-In recent years, some tracking methods based on transformer [36] have also attracted considerable attention.
-However, these methods [33, 22, 45, 41, 11, 44, 12] face a significant challenge: suboptimal inference speed.
-There are various reasons for this, such as the slow performance of the MOTR series, which relies on query-based tracking.
-The main reason is not the network structure, but issues related to hardware acceleration caused by the variable number of queries (detailed justification can be found in the appendix).
-Both our experiments and other works [24] have also demonstrated that the inference speed after lightening the network is still not very ideal.
-Regarding another recent method, MOTIP [12], while it addresses many of MOTR’s issues, its main problem lies in its complex network structure.
-Even after replacing MOTIP’s Deformable DETR [53] with RT-DETR [50], its inference speed remains unsatisfactory (results shown in Figure 1).
-And, during its implementation, some of the data is dynamic, which makes it relatively difficult to accelerate using TensorRT.
-Methods following the tracking-by-detection paradigm often claim to be extremely fast, but in reality, aside from those purely based on kinematics [4], the speed of most MOT systems is not ideal once the time required for the detection model is considered.
-Purely kinematic models typically perform worse than when they are combined with reid networks.
-So, is there a way to achieve both fast and accurate tracking?
-To address this question, we must look back at the JDT (Joint Detection and Tracking) paradigm, which was initially renowned for its speed.
-Although it has fallen out of favor in recent years due to various reasons, it likely represents a solution to the aforementioned issues.
-Previously, there was a transformer-based JDT model, known as the TransTrack [33], which shares the same pain points as MOTIP, including having two decoders.
-Other methods [22, 40, 3], similar to MOTR, track based on queries.
-CNN-based JDT models [47, 37], on the other hand, typically produce detection results and their corresponding id embeddings for association matching, with their main issue being that the models are outdated, hence they do not keep pace with current standards.
-Based on the above discussion, we present FastTrackTr, a Fast multi-object Tracking method grounded in the Transformer architecture and aligned with the JDT paradigm.
+[…]
+Multi-object tracking is a critical research field in computer vision and industrial intelligence [1, 2, 3, 4, 5], with broad applications in security monitoring, small unmanned ground vehicles, video object analysis, crowd flow detection, and other domains.
+In recent years, some tracking methodologies based on transformer [6] have garnered significant attention.
+Meanwhile, these models [7, 8, 9, 10, 11, 12, 13] have also achieved state-of-the-art (SOTA) results on many benchmarks.
+However, these methods still suffer from a critical issue: their inference speeds remain suboptimal, and the network architectures of some models are highly inconvenient for engineering deployment, making them difficult to apply in practical industrial production.
+In contrast, some other existing methods [14, 15, 16] exhibit accuracy that may lag behind Transformer-based approaches.
+Therefore, developing Transformer architectures for real-time tracking remains an urgent research goal.
+The limited inference speed of existing methods stems from multiple factors.
+For query-based tracking models like the MOTR series [9, 10, 11, 12], the primary bottleneck lies not in network architecture design but in hardware acceleration challenges caused by variable query quantities.
+Notably, computational complexity in many modules scales quadratically with the number of queries.
+Both our empirical studies and other work [17] confirm that network lightweighting alone yields insufficient speed improvements.
+The more recent method, MOTIP [13] addresses several MOTR limitations but introduces an auxiliary ID prediction network.
+Even when substituting MOTIP’s Deformable DETR [18] with RT-DETR [19], inference latency remains unsatisfactory (results shown in Fig.
+1).
+Furthermore, MOTIP’s dynamic parameterization complicates acceleration through TensorRT optimization.
+Tracking-by-Detection (TBD) paradigms [14, 15], while computationally efficient and deployment-friendly, face their own constraints.
+When accounting for the combined latency of detection models and re-identification (ReID) networks, only purely kinematic methods [16] achieve competitive speeds—at the cost of inferior performance compared to ReID-enhanced approaches.
+This raises a fundamental question: Can we develop a framework that simultaneously achieves speed, precision, and practical deployability?
+Revisiting historical solutions provides critical insights.
+The Joint Detection and Tracking (JDT) paradigm was originally designed for rapid tracking but has fallen out of favor in recent years.
+Nevertheless, its core principles may hold the key to resolving current challenges.
+Existing transformer-based JDT implementations like TransTrack [7] suffer from limitations similar to MOTIP, while other query-based methods [8, 20, 21] inherit MOTR’s fundamental constraints.
+CNN-based JDT methods [22, 23], though now less prevalent due to outdated base architectures, demonstrated an effective design: simultaneous detection output and ID embedding generation, followed by association matching.
+This paradigm inherently preserves inference efficiency while enabling deep coupling of detection features and identity representations through end-to-end joint training.
+Crucially, JDT aligns naturally with DETR architectures—the object queries produced by DETR inherently encode spatial-semantic information, thereby circumventing the tracking-by-query limitations when repurposed as ID embeddings.
+Moreover, transformer attention mechanisms surpass CNN counterparts in processing implicit temporal correlations across video frames, potentially enhancing tracking accuracy.
+Figure 1: A comparison of the frames per second (FPS) and HOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA RTX 4090.
+$\text{Dec}^{3}$ refers to a 3-layer decoder, while R18 refers to ResNet18.
+Additionally, RT-DETR was utilized to construct the MOTIP.
+The data precision of the scatter plot on the right is PyTorch fp16 and the area of points represents MOTA.
+Building upon these insights, Based on the above discussion, we present FastTrackTr, a Fast multi-object Tracking method grounded in the Transformer architecture and aligned with the JDT paradigm.
 This tracking method achieves impressive tracking progress through a unique decoder that integrates past information.
-Additionally, due to the ease of generating masks for this method, the batch size can be set relatively large, making our model considerably more training-friendly compared to some of the latest transformer-based MOT methods.
 Additionally, due to the ease of generating masks for this method, the batch size can be set relatively large, making our model considerably more training-friendly compared to some of the latest transformer-based MOT methods.
@@
   • 3)
-    Extensive experiments on multiple benchmark datasets, including DanceTrack [34], SportsMOT [7] and MOT17 [23], demonstrate that our model not only achieves competitive accuracy SOTA methods, but also significantly improves inference speed.
+    conducting extensive experiments on multiple benchmark datasets, including DanceTrack [24], SportsMOT [25], MOT17 [26], Visdrone2019 [27], and BDD100k [28], to demonstrate its robust adaptability to various environmental conditions.
+  • 4)
+    performing physical experiments in industrial settings to confirm its practical feasibility for real manufacturing applications.
 2 Related work
@@
 The initial SORT algorithm simply utilized the motion state of detected objects for association.
-Subsequent models like DeepSORT [38] and ByteTrack [48] improved tracking accuracy by enhancing detection precision and employing embedding-based Re-identification (Re-id) networks.
+Subsequent models like ByteTrack [31] improved tracking accuracy by enhancing detection precision and employing embedding-based Re-identification (Re-id) networks.
 Some approaches, like OC-SORT [16] and HybridSORT [15], improved upon SORT using purely kinematics-based methods.
-Later models such as MotionTrack [26], DeepOC-SORT [21], and SparseTrack [18] demonstrated that simultaneous improvements in both Re-id network utilization and kinematic models yield the most significant performance enhancements.
+Later models such as DeepOC-SORT [14] demonstrated that simultaneous improvements in both Re-id network utilization and kinematic models yield the most significant performance enhancements.
 Currently, there are also algorithms that directly use networks for association.
-Besides the aforementioned PuTR [16], SUSHI [6] is a tracking method based on graph neural networks.
 2.3 Joint Detection and Tracking
@@
 Here, we briefly review its development history.
-D&T [9] first proposed a multi-task architecture based on frame-based object detection and cross-frame trajectory regression.
-Subsequently, Integrated-Detection [54] improved detection performance by combining detection bounding boxes in the current frame with trajectories from the previous frame.
-Tracktor [32] directly used tracking boxes from the previous frame as candidate regions, then applied bounding box regression to provide tracking boxes for the current step, thus eliminating the box association process.
 JDE [23] and FairMOT [22] learned object detection and appearance embedding tasks from a shared backbone network.
 CenterTrack [32] localized targets through tracking-conditioned detection and predicted their offsets from the previous frame.
+MMTrack[5] has constructed a JDT model suitable for multiple scenarios.
 TransTrack used two decoders to handle tracking and detection separately, then matched and associated the results from both decoders.
@@
 Subsequently, from the second frame onwards, the historical queries from previous frames and their corresponding masks undergo processing through a single-layer encoder.
-The output and image feature of this frame are then fed into a historical-cross decoder.
-Based on the head’s output, detection results and ID embeddings are generated.
-These are then matched and associated with previous results to produce the final tracking outcome.
-This architecture achieves the efficient tracking, which can be run in real time on some devices.
-The overall structure of FastTrackTr is illustrated in Figure 2.
+The overall structure of FastTrackTr is illustrated in Fig.
+2.
 In frame $T_{1}$, a standard decoder is used to initialize the initial historical information.
@@
 In the next frame, a Historical Decoder, which includes a historical cross-attention layer compared to normal decoder layers, processes this historical information.
-The tracking principle is very similar to that of JDT and FairMOT, utilizing an additional ID Embedding head to obtain the object’s appearance embedding, which is then used along with detection results for association and matching.
-It is worth noting that the backbone network is ResNet [14].
+The tracking principle is very similar to that of JDE and FairMOT, utilizing an additional ID Embedding head to obtain the object’s appearance embedding, which is then used along with detection results for association and matching.
+This architecture achieves efficient tracking, which can be run in real time on edge devices.
 3.2 Learning Detection and ID Embedding
-To enable our FastTrackTr to simultaneously output both the target’s position and appearance embedding in a single forward pass, we formulate the problem as follows: Suppose we have a training dataset ${\mathbf{I},\mathbf{B},\mathbf{y}}_{i=1}^{N}$.
-Here, $\mathbf{I}\in\mathbb{R}^{c\times h\times w}$ represents an image frame, and $\mathbf{B}\in\mathbb{R}^{k\times 4}$ denotes the bounding box annotations for the $k$ targets in this frame.
-$\mathbf{y}\in\mathbb{Z}^{k}$ represents the partially annotated identity labels, where -1 indicates targets without an identity label.
-We hope model can output predicted bounding boxes $\hat{\mathbf{B}}\in\mathbb{R}^{\hat{k}\times 4}$ and appearance embeddings $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{k}\times D}$, where $D$ is the dimension of the embedding.
-For the output box, this requirement is easily met, as DETR[5] itself is a detection model.
-To achieve rapid performance, this paper employs RT-DETR [50] as the base model, which inherently possesses state-of-the-art detection capabilities.
-We have not made significant modifications to its detection component.
-The loss function for this part remains consistent with RT-DETR, as shown below:
-[…]
-where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}=$$\{\hat{c},\hat{b}\}$ and $y=\{c,b\},c$ and $b$ represent categories and bounding boxes, respectively.
-Similar to RT-DETR, we incorporate the IoU score into the objective function of the classification branch, following the approach from VFL [46], to enforce consistency between the classification and localization of positive samples.
-$\mathcal{L}_{box}$ includes the L1 loss and the generalized IoU loss [27], which are not described here.
-And, the auxiliary loss from DAB-DETR [17] is also employed in our approach.
-The second objective addresses a metric learning problem, specifically the learning of an embedding space where instances of the same identity are close to each other while those of different identities are far apart.
-Previous JDT methods (e.g., JDE [37], FairMOT [47]) rely on one-hot labels to construct ID embeddings, necessitating predefined ID quantities to configure embedding dimensions.
-This design introduces two critical limitations: (1) explosive dimensionality growth in dense scenes and (2) severely restricted generalization due to dataset-specific ID dependencies.
-Subsequent Transformer-based trackers inherit these constraints through their query mechanisms.
-While recent approaches like MOTIP [12] (using ID dictionaries) and PuTR (leveraging embedding similarity) improve generalizability, their trajectory decoders overemphasize motion patterns, conflicting with FastTrackTr’s queries that encode richer spatial features and positional priors.
-This mismatch hinders effective convergence.
-To resolve these issues, we reformulate ID embedding learning as a re-identification (ReID) problem, decoupling embedding dimensions from ID quantities and prioritizing feature distance/similarity optimization.
-This paradigm shift enhances cross-scene adaptability and better aligns with image-driven tracking frameworks.
-For the above reasons and for better performance, in this paper, we adopt the more advanced circle loss [35] in this work.
-Specifically, by utilizing the dot product similarity, the id embeddings of the same object across multiple frames are regarded as positive pairs indicated as $s_{p}$, while those between distinct objects are treated as negative samples represented by $s_{n}$.
+To enable our FastTrackTr to simultaneously output both the target’s position and id embedding in a single forward pass, ensuring optimal inference speed, we formulate the problem as follows: Given a training dataset $\{(\mathbf{I}_{i},\mathbf{B}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ consisting of image frames $\mathbf{I}_{i}\in\mathbb{R}^{C\times H\times W}$, bounding box annotations $\mathbf{B}_{i}\in\mathbb{R}^{K_{i}\times 4}$, and partial identity labels $\mathbf{y}_{i}\in\mathbb{Z}^{K_{i}}$, we aim to develop a unified model that simultaneously predicts bounding boxes $\tilde{\mathbf{B}}_{i}\in\mathbb{R}^{\tilde{K}_{i}\times 4}$ and discriminative appearance embeddings $\tilde{\mathbf{F}}_{i}\in\mathbb{R}^{\tilde{K}_{i}\times D}$ through a single forward pass.
+The model must satisfy two fundamental requirements: (1) accurate bounding box regression through minimization of the localization loss $\mathcal{L}_{\text{det}}(\tilde{\mathbf{B}},\mathbf{B},\tilde{\mathbf{y}},\mathbf{y})$, and (2) temporally consistent feature embeddings where the distance $d_{\phi}(f_{t}^{k},f_{t+\Delta t}^{m})$ between matching identities ($\mathbf{y}_{t}^{k}=\mathbf{y}_{t+\Delta t}^{m}$) is smaller than non-matching pairs ($\mathbf{y}_{t}^{k}\neq\mathbf{y}_{t+\Delta t}^{m^{\prime}}$) by a margin $\alpha>0$.
+Formally, the joint objective can be expressed as:
+$\min_{\theta,\phi}\mathbb{E}\left[\mathcal{L}_{\text{det}}+\lambda\max(0,d_{\phi}(f_{t}^{k},f_{t+\Delta t}^{m})-d_{\phi}(f_{t}^{k},f_{t+\Delta t}^{m^{\prime}}))\right]$ (1)
+where $\lambda$ controls the balance between detection accuracy and feature discriminability, with $f_{t}^{k}$ denoting the $k$-th row vector of $\tilde{\mathbf{F}}_{t}$ corresponding to the $k$-th predicted instance.
+Our framework builds upon the DETR architecture [18], selected for its inherent compatibility with our technical objectives.
+The Transformer decoder’s parallel decoding mechanism transforms learnable object queries into interpretable target representations through cross-attention interactions with image features.
+These joint representations intrinsically encode spatial coordinates and semantic features, which are conventionally mapped to classification probabilities and bounding box coordinates via separate prediction heads.
+We extend this paradigm through two strategic modifications: (1) Introducing an auxiliary ID embedding head implemented as a fully-connected layer that projects object queries into a discriminative feature space for target association, and (2) Implementing a joint optimization strategy that preserves detection performance while learning identity-aware representations.
+The base detection component employs RT-DETR [19] for its real-time efficiency, maintaining its original dual-branch structure and loss formulation:
+$\mathcal{L}_{\text{det}}=\underbrace{\mathcal{L}_{\text{box}}(\hat{b},b)}_{\text{Localization}}+\underbrace{\mathcal{L}_{\text{cls}}(\hat{c},c,\text{IoU})}_{\text{Classification}}$ (2)
+where $\hat{b},b$ denote predicted/ground-truth boxes and $\hat{c},c$ represent class predictions.
+The classification loss integrates IoU-aware weighting following RT-DETR, while the box regression combines L1 and generalized IoU losses.
+Previous JDT methods [23, 22] suffer from scalability limitations due to their one-hot identity encoding schemes.
+We instead formulate identity embedding learning as a metric learning problem, decoupling feature dimensions from identity quantities to enhance cross-scene generalizability.
+This contrasts with recent motion-centric approaches [13, 30] that underutilize spatial features encoded in DETR queries.
+Our method employs Circle Loss [33] with adaptive margin control to optimize pairwise similarities in the embedding space.
+Specifically, by utilizing the dot-product similarity, the id embeddings of the same object across multiple frames are regarded as positive pairs indicated as $s_{p}$, while those between distinct objects are treated as negative samples represented by $s_{n}$.
 - $\alpha_{p}$ and $\alpha_{n}$ are the adaptive weighting factors for positive and negative samples, determined by the following equations:
-$\alpha_{p}=\left[O_{p}-s_{p}\right]^{+},\quad\alpha_{n}=\left[s_{n}-O_{n}\right]^{+}$ (2)
-where $O_{p}=1+m$ and $O_{n}=-m$, with $m$ being a margin hyperparameter.
-These terms adjust the weight of each sample based on its distance from the ideal similarity value.
-The margins $\Delta_{p}=1-m$ and $\Delta_{n}=m$ are used to further separate positive and negative pairs in the embedding space.
-The logits for positive and negative pairs are then computed as:
-$\log_{p}=-\gamma\alpha_{p}(s_{p}-\Delta_{p}),\quad\log_{n}=\gamma\alpha_{n}(s_{n}-\Delta_{n})$ (3)
-Finally, the unified Circle Loss function is formulated as follows:
-$&\mathcal{L}_{\text{reid}}=\\ &\log\left[1+\sum_{j=1}^{L}\exp\left(\gamma\left(\alpha_{n}^{j}\left(s_{n}^{j}-\Delta_{n}\right)\right)\right)\sum_{i=1}^{K}\exp\left(-\gamma\left(\alpha_{p}^{i}\left(s_{p}^{i}-\Delta_{p}\right)\right)\right)\right]\end{aligned}$
-This formulation allows for a dynamic weighting of positive and negative pairs, enhancing the model’s ability to discriminate between similar and dissimilar identities.
-Circle Loss has proven to be an effective solution for embedding learning in tasks that require fine-grained discrimination, which aligns with the goals of appearance embedding in our tracking framework.
-The final total loss is as follows:
-$\mathcal{L}_{overall}=\lambda_{reid}*\mathcal{L}_{reid}+\lambda_{det}*\mathcal{L}_{det}$ (5)
-where $\lambda_{reid}$ and $\lambda_{det}$ are proportional coefficients used to balance the loss.
-In this paper, both of them are set to 1.
-Notably, it should be noted that, when computing the ReID loss, similar to the detection loss, the outputs from each decoder layer are incorporated into the calculation.
+$\alpha_{p}$ $=\max(0,O_{p}-s_{p}),\quad\alpha_{n}=\max(0,s_{n}-O_{n})$ (3)
+$O_{p}$ $=1+m,\quad O_{n}=-m\quad(m>0)$
+where $m$ controls the separation margin. The final loss formulation:
+$\mathcal{L}_{\text{reid}}=\log\left[1+\sum_{j}e^{\gamma\alpha_{n}(s_{n}^{j}-\Delta_{n})}\sum_{i}e^{-\gamma\alpha_{p}(s_{p}^{i}-\Delta_{p})}\right]$ (4)
+with $\Delta_{p}=1-m$, $\Delta_{n}=m$, and $\gamma$ as the scaling factor. This adaptive weighting mechanism prioritizes hard samples in both positive and negative groups.
+The complete objective function combines detection and metric learning losses through layer-wise aggregation across decoder stages:
+$\mathcal{L}_{\text{total}}=\lambda_{\text{det}}\mathcal{L}_{\text{det}}+\lambda_{\text{reid}}\mathcal{L}_{\text{reid}}$ (5)
+where $\lambda_{\text{det}}=\lambda_{\text{reid}}=1$ ensures balanced optimization.
+At the same time, in order to better enable the model to implicitly learn temporal information, like other Transformer-based MOT models [9, 13, 8, 7], the loss here is updated after accumulating several consecutive frames.
 3.3 Historical Decoder and Encoder
+While preceding sections have detailed the construction of an efficient Transformer-based MOT framework, the inherent temporal aggregation capabilities of Transformers remain underexploited in our current implementation.
+This section introduces a systematic methodology for efficiently leveraging historical information within the current frame.
 Historical Decoder: Previous DETR-based MOT models [9, 8, 11] typically feed the final output queries of the previous frame, after some processing, directly into the decoder of the current frame to achieve tracking.
@@
 This significantly reduces the computational load.
-The differences between our decoder and the standard decoder are shown in Figure 3.
+The differences between our decoder and the standard decoder are shown in Fig.
+3.
 Specifically, each decoder module in DETR [18] consists of a sequential structure comprising a self-attention layer for the queries, a visual cross-attention layer, and a feed-forward neural network (FFN).
@@
 The queries with estimated historical information are then input into the inter-query self-attention layer for feature interaction between objects, and the visual cross-attention layer for collecting visual semantics.
-Figure 3: Schematic of the flow of our historical-cross decoder and a general decoder. Compared to a general decoder, ours replaces self-attention with historical cross-attention.
+Figure 3: Schematic of the flow of our historical-cross decoder and a general decoder.
+Compared to a general decoder, ours replaces self-attention with historical cross-attention.
+In our historical cross-attention mechanism, the setting of $q$ remains the same, while $k$ and $v$ incorporate information from historical frames.
 Historical Encoder: In FastTrackTr, we have introduced an additional Historical Encoder to enhance the modeling of temporal relationships, and we utilize a masking mechanism during the self-attention process to provide contextual priors for the tracked objects.
@@
 During training, the generation of masks is determined by the ground truth labels, while in inference, masks are based on confidence scores to decide which objects to retain.
-The workflow of the historical encoder is illustrated in Figure 4.
+The workflow of the historical encoder is illustrated in Fig.
+4.
 After generating the mask, we fuse the decoder output $q_{f}^{t}$ with the historical information $q_{f}^{t-1^{\prime}}$ through an attention mechanism, computed as follows:
 $Q_{E}=K_{E}=\operatorname{Linear}(q^{t}_{f}+q_{pos}),V_{E}=\operatorname{Linear}(q^{t-1^{\prime}}_{f})$ (8)
-where $q_{pos}$ represents a set of learnable positional encodings and the attention weights are computed via scaled dot-product as previously used in the decoder, followed by masking to suppress less relevant historical elements.
+where $q_{pos}$ represents positional encodings and the attention weights are computed via scaled dot-product as previously used in the decoder, followed by masking to suppress less relevant historical elements.
 Finally, the fused features are then processed by a feed-forward network to obtain $q^{t^{\prime}}_{f}$.
@@
 where $\tilde{f}$ indicates the embedding of the assigned observation, and $f_{t}$ indicates the embedding of the tracklet at frame $t$, $\eta$ is a momentum term for smoothing, and we set $\eta=0.9$.
-Of course, considering that there are now many optimized methods for this process, we have also implemented several validated, simple, and effective techniques to enhance this workflow.
-Specific improvements and their impacts can be found in the appendix.
+Meanwhile, to further improve the model’s performance, similar to MMTrack [5], we adopted the iterative strategy of BYTE [31] and ID features, which is consistent with the proposals of [22, 16].
 4 Experiments
@@
 Similarly, the latency of our method accounts for its associated time consumption, and these are highlighted in color.
-$\text{Dec}^{3}$ refers to a 3-layer decoder, while R18 refers to ResNet18.
-Unless otherwise specified, all Transformer-based models have ResNet50 for Backbone and 6 layers for decoder.
-Methods Image Res. Numerical Precision Overall Latency(ms) HOTA Overall Latency with TensorRT FP16(ms)
-Deep OC-SORT*[21] 1333×800 FP16 36.6ms (27.8fps) 58.5 13.5ms(74.1fps)
-SUSHI*[6] 1333×800 FP32/FP16 42.7ms (23.4fps) / 38.0ms (26.3fps) 58.7 -
-PuTR*[16] 1333×800 FP32/FP16 29.9ms (33.4fps) / 27.7ms (36.1fps) 54.5 -
+Unless otherwise specified, all Transformer-based models have ResNet50 for Backbone and 6 layers for decoder besides PuTR as it is a method of TBD paradigm.
+Not providing methods to accelerate TensorRT implies that they may not be well-suited for it.
+Methods Image Res. Numerical Precision Overall Latency(ms) HOTA
+Overall Latency with
+TensorRT FP16(ms)
+DiffMOT*[35] 1333×800 FP16 43.3ms (23.1fps) 55.7 22.3ms(44.8fps)
+PuTR*[30] 1333×800 FP32/FP16 29.9ms (33.4fps) / 27.7ms (36.1fps) 54.5 15.8ms(63.3fps)
 MOTR[9] 1333×800 FP32/FP16 60.0ms (16.5fps) / 52.9ms (18.9fps) 51.3 -
 MOTIP[13] 1333×800 FP32/FP16 61.7ms (16.2fps) / 45.7ms (21.8fps) 59.1 -
-FastTrackTr 1333×800 FP32/FP16 40.2ms (24.8fps) / 30.9ms (32.4fps) 56.8 11.6ms(86.6fps)
+FastTrackTr 1333×800 FP32/FP16 40.2ms (24.8fps) / 30.9ms (32.4fps) 56.9 11.6ms(86.6fps)
 MOTIP(RT-DETR) 640×640 FP32/FP16 32.1ms (31.2fps) / 24.5ms (40.8fps) 56.2 -
-FastTrackTr 640×640 FP32/FP16 23.6ms (42.3fps) / 19.6ms (51.0fps) 54.1 6.0ms(166.4fps)
+FastTrackTr 640×640 FP32/FP16 23.6ms (42.3fps) / 19.6ms (51.0fps) 54.8 6.0ms(166.4fps)
 FastTrackTr-Dec3 640×640 FP32/FP16 20.4ms (49.0fps) / 14.8ms (67.6fps) 51.2 4.9ms(203.7fps)
@@
 4.1 Setting
-Datasets: Our proposed method mainly evaluated on DanceTrack [34] and SportsMOT [7], two recently introduced datasets with large-scale training data, which helps prevent overfitting.
-Both datasets also provide official validation sets, allowing for exploratory experiments.
-Additionally, we present results on the MOT17 [23] benchmark.
+Datasets: To evaluate the tracking performance of FastTrackTr in real-world scenarios, we selected five datasets for testing: Dancetrack [24], SportMOT [25], MOT17 [26], BDD100k [28], and VisDrone2019 [27].
+These datasets comprehensively encompass most real-world scenarios—for instance, MOT17 for surveillance, BDD100k for driving, VisDrone2019 for drone-based applications, and DanceTrack/SportMOT for sports-related contexts.
+Finally, we also tested our model on our own physical platform.
 Metrics: We assessed our method using widely recognized MOT evaluation metrics.
@@
 These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance.
-Implementation Details: For inference speed, our FastTrackTr is built on RT-DETR [50], with ResNet50 [14] as the backbone unless specified otherwise.
-The model is implemented in PyTorch [25], and the training is conducted on 8 NVIDIA RTX 4090 GPUs.
+We also present additional evaluation metrics, including false positives (FP), false negatives (FN), identity switches (IDS), detection accuracy (DetA), association accuracy (AssA), and frame rate (FPS).
+And, for BDD100K, to better evaluate the performance of multiclass and multiobject tracking, we use tracking accuracy of all things (TETA) [28], localization accuracy (LocA), association accuracy (AssocA) and classification accuracy (ClsA) metrics.
+Implementation Details: For inference speed, our FastTrackTr is built on RT-DETR [19], with ResNet50 as the backbone unless specified otherwise.
+The model is implemented in PyTorch, and the training is conducted on 8 NVIDIA RTX 4090 GPUs.
 The sequence length for training is set to 8 unless otherwise noted.
-All ablation experiments use an image size of $640\times 640$.
-For further testing of the best tracking accuracy of our method, we use an image size of $800\times 1333$ when comparing with SOTA methods.
+All ablation experiments use an image size of $\times 640$ 640.
+For further testing of the best tracking accuracy of our method, we use an image size of 800×1333800$\times$ 1333800 × 1333 when comparing with SOTA methods.
 The Historical Encoder during training applies random mask transformations, with the probability related to the number of ground-truth objects.
@@
 For all three datasets, $n_{q}$ is set to 300.
-Dancetrack and SportMOT are trained for 27 epochs, while MOT17 is trained for 80 epochs.
-The optimizer is AdwmW [19] with a learning rate of $10^{-4}$ and weight decay of $5.0\times 10^{-4}$.
-Due to GPU memory limitations, the batch size is set to 4 for smaller images and 1 for larger images.
+Dancetrack, BDD100k and SportMOT are trained for 27 epochs, while MOT17 and VisDrone2019 are trained for 80 epochs.
+The optimizer is AdwmW with a learning rate of $10^{-4}$ and weight decay of $5.0\times 10^{-4}$.
+Due to GPU memory limitations, the batch size is set to 4 for smaller images and 1 for larger images on one GPU.
 Lastly, for all datasets, the final output dimension of ID embedding is 256.
@@
 Within this benchmark, all models undergo acceleration via TensorRT, with computational latency measurements exclusively encompassing inference and post-processing durations.
-This paper adopts the same configuration while substituting the detection dataset COCO2017 [15] with the MOT dataset Dancetrack .
-However, this substitution reveals a critical implementation challenge: most transformer-based MOT methods [33, 22, 45, 41, 11, 44, 12, 16] developed in PyTorch incorporate dynamic computational graphs during inference, presumably to reduce computational overhead or simplify implementation.
+This paper adopts the same configuration while substituting the detection dataset with the MOT dataset Dancetrack .
+However, this substitution reveals a critical implementation challenge: many transformer-based MOT methods [7, 8, 9, 10, 11, 12, 13] developed in PyTorch incorporate dynamic computational graphs during inference, presumably to reduce computational overhead or simplify implementation.
 These dynamic elements create significant compatibility issues (specifically non-fixed tensor shapes and conditional branching) when converting models to TensorRT - challenges that remain unresolved in our current implementation.
 To ensure fair comparison, Table 1 comprehensively presents both speed and accuracy metrics across tracking methods using various numerical precisions in PyTorch.
-Notably, for our proposed models and the TBD method, only the network benefits from TensorRT acceleration, while the association module retains its original Python implementation.
+Notably, for FastTrackTr and the TBD method, only the network benefits from TensorRT acceleration, while the association module retains its original Python implementation and similar to real-time detection, in this paper, we only measure the time consumption of the tracking system, and the time for data preprocessing and post-processing is not included.
 Table 2: Performance comparison between FastTrackTr and existing methods on the Dancetrack [24] test set.
@@
 FairMOT[22] 39.7 66.7 23.8 82.2 40.8
-CenterTrack[51] 41.8 78.1 22.6 86.8 35.7
-TraDeS[39] 43.3 74.5 25.4 86.2 41.2
 TransTrack[7] 45.5 75.9 27.5 88.4 45.2
+MMTrack[5] 46.5 69.9 31.0 86.1 48.3
 ByteTrack[31] 47.7 71.0 32.1 89.6 53.9
-GTR[52] 48.0 72.5 31.9 84.7 50.3
 QDTrack[38] 54.2 80.1 36.8 87.7 50.4
@@
 OC-SORT[16] 55.1 80.3 38.3 92.0 54.6
-MoveSORT[1] 56.1 81.6 38.7 91.8 56.0
 PuTR[30] 55.8 /// /// 91.9 58.2
-C-BIoU[42] 60.6 81.3 45.4 91.6 61.6
-CMTrack[31] 61.8 /// 46.4 92.5 63.3
+Deep OC-SORT[14] 61.3 82.2 45.8 92.3 61.5
+DiffMOT[35] 62.3 82.2 47.2 92.8 63.0
 FastTrackTr(ours) 62.4 78.4 49.8 88.8 64.8
 Comparison of State-Of-The-Art: To better demonstrate the superiority of our FastTrackTr, we employed nearly all available data augmentations and used larger image sizes.
-Tables 2, 4, and 3 respectively showcase the maximum tracking performance of FastTrackTr on DanceTrack, SportsMOT, and MOT17 benchmark.
+Tables 2, 3 ,4, 5 and 6 respectively showcase the maximum tracking performance of FastTrackTr on all benchmarks.
 Similar to many previous works [32, 29, 11, 9], we utilized the CrowdHuman [39] dataset for data augmentation on MOT17 and DanceTrack.
@@
 This indicates that our method performs better even when the dataset images do not allow for sufficient training of the model, suggesting that our approach has potential for practical application in real-world settings, rather than being confined mostly to academic research like previous transformer methods, with few exceptions.
+And, We also provide various datasets and visualization results verified on physical objects, as shown in Fig.
+5.
 Table 3: Performance comparison with state-of-the-art methods on MOT17 [26]. The best performance among the transformer-based methods is marked in bold.
@@
 CNN based:
-CenterTrack [51] 52.2 53.8 51.0 67.8 64.7
-QDTrack [10] 53.9 55.6 52.7 68.7 66.3
 FairMOT [22] 59.3 60.9 58.0 73.7 72.3
+MMTrack[5] 61.1 - - 74.9 74.4
 ByteTrack [31] 63.1 64.5 62.0 80.3 77.3
 OC-SORT [16] 63.2 63.2 63.4 78.0 77.5
-C-BIoU [42] 64.1 64.8 63.7 81.1 79.7
-MotionTrack [26] 65.1 65.4 65.1 81.1 80.1
+DiffMOT [35] 64.5 65.7 64.6 79.8 79.3
 Transformer based:
@@
 FastTrackTr(ours) 70.1 83.9 58.6 94.0 72.0
+Table 5: Comparison With the SOTA Methods on the VisDrone2019 Test Set
+Methods MOTA $\uparrow$ IDF1 $\uparrow$ FP $\downarrow$ FN $\downarrow$ IDS $\downarrow$
+MOTR [9] 22.8 41.4 28407 147937 959
+TrackFormer [8] 25.0 30.5 25856 141526 4840
+MMTrack [5] 36.7 54. 7 23849 120839 545
+FastTrackTr 42.8 62.4 16474 85367 309
+Table 6: Comparison to existing methods on the BDD100K validation set. Best results are marked in bold.
+Methods TETA LocA AssocA ClsA
+QDTrack [38] 47.8 45.8 48.5 49.2
+MOTRv2 [29] 54.9 49.5 51.9 63.1
+CO-MOT [10] 52.8 48.7 56 2 63.6
+FastTrackTr 55.1 49.8 54 2 65.3
+Figure 5: The visualization of FastTrackTr on various datasets. Ours represents the actual measured data of our driverless car on the road, and the model used here is trained on BDD100k.
 4.3 Ablation Study
@@
 The ablation study in Table 7 highlights the progressive enhancements of FastTrackTr on the DanceTrack validation set.
-Starting with a baseline HOTA of 47.3, adding the Decoder increases HOTA to 51.6, AssA to 37.2, and IDF1 to 51.3.
-Incorporating both Decoder and Encoder further boosts HOTA to 53.8, AssA to 39.9, and IDF1 to 53.5.
-Finally, including the Mask module achieves the highest scores with HOTA of 54.1, AssA of 40.0, and IDF1 of 54.7.
-These results demonstrate that each component significantly contributes to the overall tracking performance of FastTrackTr.
+Starting with a baseline HOTA of 47.3, adding the Decoder increases HOTA to 51.6.
+Incorporating both Decoder and Encoder further boosts HOTA to 53.8.
+Finally, including the Mask module achieves the highest scores with HOTA of 54.1.
+In terms of the improvement degree, the Decoder with the introduction of historical information brings about the most significant performance enhancement.
+Adding an Encoder that further aggregates the temporal information of historical frames also contributes to a performance increase.
+The introduction of mask not only shields off useless historical information but also the randomness generated during its training can boost the generalization ability of the model.
 Table 7: Ablation Study of our proposed FastTrackTr on the DanceTrack validation set.
@@
 It is important to note that in this table, $q_{f}^{t-1}$ has not passed through the historical encoder, hence its performance is not as good as that of the final model.
+The advantage of Concat lies in combining the features of the current query and historical states.
+By leveraging the complementarity of spatiotemporal information and the dynamic balancing of the attention mechanism, it significantly boosts the joint performance of detection and tracking.
 Table 8: The Impact of Different Historical Cross Attention mechanisms on Tracking Performance
@@
 concat concat 51.6 72.8 37.2 51.3
-Table 7 analyzes the effect of different length of video clips during training.
-Utilizing 8 frames yields the highest HOTA, demonstrating superior tracking accuracy.
-Compared to using fewer frames, 8 frames offer substantial performance improvements while maintaining efficient training times.
-Beyond 8 frames, metrics show minimal or negligible gains, indicating diminishing returns with additional frames.
-Thus, selecting 8 training frames provides an optimal balance between enhanced performance and training efficiency for FastTrackTr.
-Table 7: The length of video clips during training.
-Training frames HOTA DetA AssA IDF1
-3 51.7 72.8 36.7 50.9
-5 53.5 74.2 38.8 53.5
-8 54.1 75.6 40.0 54.7
-9 54.0 75.4 39.6 54.8
-10 54.1 75.8 39.8 54.6
 In Table 9, we demonstrate the impact of using circle loss [33], triplet loss [40], and one-hot encoding loss of the JDE [23] on our method performance during the training of ID embeddings.
@@
 In comparison, triplet loss shows the lowest performance, and one-hot encoding loss provides intermediate results.
-These results demonstrate that circle loss effectively enhances the training of ID embeddings, leading to superior tracking performance in our method.
-Table 8: The impact of various ID Embedings losses on performance.
+These results demonstrate that circle loss effectively enhances the training of ID embeddings for, leading to superior tracking performance in our method.
+Table 9: The impact of various ID Embeddings losses on performance.
 $loss$ HOTA DetA AssA IDF1
@@
 circle loss 54.1 75.6 40.0 54.7
-5 Conclusion and Limitations
-Conclusion: In this paper, we introduce a rapid tracking method, FastTrackTr, which demonstrates superior performance across multiple datasets.
+4.4 Experiments in Actual Application Scenarios
+Figure 6: Unmanned vehicle for physical verification.
+Table 10: The speed and accuracy of our method and some existing methods using TensorRT FP16 precision on edge devices on our own verification platform and the accuracy of these models on Dancetrack are provided for reference.
+Methods Image Res. Overall Latency(ms) HOTA
+Deep OC-SORT[14] 1333 $\times$ 800 163.9 ms(6.1 fps) 58.5
+PuTR [30] 1333 $\times$ 800 67.3ms(14.8fps) 54.5
+OC-SORT [16] 1333 $\times$ 800 54.3 ms(18.4 fps) 52.1
+FastTrackTr 1333 $\times$ 800 48.7 ms(20.5 fps) 56.9
+Deep OC-SORT [14] 640 $\times$ 640 141.1 ms(7.1 fps) -
+OC-SORT [16] 640 $\times$ 640 36.7 ms(27.2 fps) -
+FastTrackTr 640 $\times$ 640 35.6 ms(28.1 fps) 54.8
+FastTrackTr-Dec3 640 $\times$ 640 30.1ms(33.2fps) 51.2
+FastTrackTr-R18-Dec3 640 $\times$ 640 16.8ms(59.2fps) 47.4
+To validate the performance of our method in practical industrial environments, we conducted tests on the autonomous vehicle shown in Fig.
+6.
+The core processor of this autonomous vehicle is NVIDIA’s Jetson AGX Orin, and the vehicle is equipped with sensors such as a lidar, an IMU, a GPS module, a fisheye camera, and an RGB camera.
+A Redmi AX3000 router is used to build an in-vehicle local area network, satisfying the network communication between the host and the lidar as well as the data transmission between the client and the autonomous vehicle host.
+To further simulate real-world scenarios, in addition to the tracking method, a SLAM algorithm is also running on the autonomous vehicle.
+The final time consumption of each tracking method is shown in Table 10, and the indicators marked in the table are the accuracy of the methods on Dancetrack.
+Since the official details on how OC - SORT trains the YOLOX detector are not provided, the accuracy of our trained model is not ideal, which may cause some misunderstandings.
+Therefore, we do not provide the specific tracking accuracy of OC - SORT and Deep OC - SORT under the image size of 640×640, and only provide the speed.
+5 Conclusion
+In this paper, we introduce a real time tracking method, FastTrackTr, which demonstrates superior performance across multiple datasets.
 By incorporating novel trajectory encoder and decoder designs, we effectively enhance the accuracy and robustness of tracking.
@@
 Future work will focus on further optimizing the model structure to adapt to more complex scenarios and exploring deployment possibilities across different hardware platforms.
-Limitations: The primary limitation of FastTrackTr is its tracking accuracy, which does not surpass that of MOTIP.
-This issue likely arises because MOTIP employs a specialized network structure for ID encoding and its training method for long sequences, which ensures better temporal information learning.
+However, our method currently does not surpass the tracking accuracy of MOTIP, primarily due to its specialized network architecture for ID encoding and sequence-optimized training paradigm that captures long-term temporal dependencies.
+which ensures better temporal information learning.
 As the saying goes, ”there are trade-offs”; these aspects also result in poor inference speed and slow training for MOTIP.
 Conversely, FastTrackTr does not exhibit these drawbacks.
-Appendix A More discussion on FastTrackTr
-A.1 Description of calculations
-In this section, we analyze the computational complexity of FastTrackTr and MOTR.
-The goal is to determine the threshold at which the computational cost of the MOTR exceeds that of our FastTrackTr when the number of queries increases.
-We focus on the decoder layer, including the self-attention, cross-attention, feed-forward network (FFN), and normalization layers.
-A.1.1 Model Definitions
-MOTR (Variable Query Number): The number of queries increases from $N$ to $N_{q}=N+\Delta N$, where $\Delta N$ is the additional number of queries.
-Both the self-attention and subsequent components are affected by this increase.
-FastTrackTr (Modified Self-Attention): The self-attention mechanism is altered such that the key and value tensors have dimensions $2N\times C$, while the query tensor remains $N\times C$.
-Other components retain their original computational complexity.
-A.1.2 Notations
-  • $N$: Original number of queries.
-  • $\Delta N$: Additional queries in MOTR ($N_{q}=N+\Delta N$).
-  • $C$: Dimensionality of each query.
-  • $M$: Length of the memory (number of keys/values in cross-attention).
-  • $d_{\text{ff}}$: Dimensionality of the FFN’s hidden layer (typically $4C$).
-A.1.3 Computational Complexity Analysis
-MOTR:
-  1. 1.
-    Self-Attention:
-    $\mathcal{O}_{\text{SA}_{1}}$ $=(N+\Delta N)^{2}\cdot C$ (10)
-  2. 2.
-    Cross-Attention:
-    $\mathcal{O}_{\text{CA}_{1}}$ $=(N+\Delta N)\cdot M\cdot C$ (11)
-  3. 3.
-    Feed-Forward Network (FFN):
-    $\mathcal{O}_{\text{FFN}_{1}}$ $=2(N+\Delta N)\cdot C\cdot d_{\text{ff}}$ (12)
-    (Two linear layers with activation in between.)
-  4. 4.
-    Layer Normalization:
-    $\mathcal{O}_{\text{LN}_{1}}$ $=2(N+\Delta N)\cdot C$ (13)
-    (Two LayerNorm layers.)
-Total Complexity of MOTR:
-$\mathcal{O}_{1}$ $=(N+\Delta N)^{2}\cdot C+(N+\Delta N)\cdot M\cdot C$
-$\quad+2(N+\Delta N)\cdot C\cdot d_{\text{ff}}+2(N+\Delta N)\cdot C$ (14)
-FastTrackTr:
-  1. 1.
-    Self-Attention:
-    $\mathcal{O}_{\text{SA}_{2}}$ $=N\cdot(2N)\cdot C=2N^{2}\cdot C$ (15)
-  2. 2.
-    Cross-Attention:
-    $\mathcal{O}_{\text{CA}_{2}}$ $=N\cdot M\cdot C$ (16)
-  3. 3.
-    Feed-Forward Network (FFN):
-    $\mathcal{O}_{\text{FFN}_{2}}$ $=2N\cdot C\cdot d_{\text{ff}}$ (17)
-  4. 4.
-    Layer Normalization:
-    $\mathcal{O}_{\text{LN}_{2}}$ $=2N\cdot C$ (18)
-Total Complexity of FastTrackTr:
-$\mathcal{O}_{2}$ $=2N^{2}\cdot C+N\cdot M\cdot C+2N\cdot C\cdot d_{\text{ff}}+2N\cdot C$ (19)
-A.1.4 Inequality for Computational Cost Comparison
-To find the threshold $\Delta N$ where $\mathcal{O}_{1}>\mathcal{O}_{2}$, we set up the inequality:
-$(N+\Delta N)^{2}\cdot C+(N+\Delta N)\cdot M\cdot C$
-$\quad+2(N+\Delta N)\cdot C\cdot d_{\text{ff}}+2(N+\Delta N)\cdot C$
-$>2N^{2}\cdot C+N\cdot M\cdot C+2N\cdot C\cdot d_{\text{ff}}+2N\cdot C$ (20)
-Simplifying the inequality by eliminating common terms and dividing both sides by $C$:
-$(N+\Delta N)^{2}-2N^{2}+(N+\Delta N-N)\cdot M$
-$\quad+2(N+\Delta N-N)\cdot d_{\text{ff}}+2(N+\Delta N-N)>0$
-$\Rightarrow$ $(N+\Delta N)^{2}-2N^{2}+\Delta N\cdot(M+2d_{\text{ff}}+2)>0$ (21)
-Expanding the square term:
-$N^{2}+2N\Delta N+(\Delta N)^{2}-2N^{2}$
-$+\Delta N\cdot(M+2d_{\text{ff}}+2)>0$ (22)
-Combining like terms:
-$-N^{2}+2N\Delta N+(\Delta N)^{2}$
-$+\Delta N\cdot(M+2d_{\text{ff}}+2)>0$ (23)
-Simplifying:
-$-N^{2}+\Delta N\left(2N+\Delta N+M+2d_{\text{ff}}+2\right)>0$ (24)
-A.1.5 Determining the Threshold $\Delta N$
-To find the minimum $\Delta N$ satisfying the inequality, we proceed as follows.
-$-N^{2}+\Delta N\left(2N+\Delta N+M+2d_{\text{ff}}+2\right)>0$ (25)
-A.1.6 Final Solution with Specific Numerical Values
-Assuming typical values for a Deformable DETR decoder layer:
-  • Number of original queries: $N=300$
-  • Memory length: $M=8400$
-  • Query dimensionality: $C=256$
-  • FFN hidden dimension: $d_{\text{ff}}=4C=1024$
-Substituting the numerical values:
-$-(300)^{2}+\Delta N\Big{(}2\times 300+\Delta N+8400$
-[…]
-Computing constants:
-$-90,000+\Delta N\left(600+\Delta N+8400+2048+2\right)>0$ (27)
-Simplifying the expression inside the parentheses:
-$-90,000+\Delta N\left(\Delta N+11,050\right)>0$ (28)
-Forming a quadratic inequality:
-$(\Delta N)^{2}+11,050\Delta N-90,000>0$ (29)
-Solving the quadratic equation:
-$(\Delta N)^{2}+11,050\Delta N-90,000=0$ (30)
-Using the quadratic formula:
-$\Delta N=\frac{-11,050\pm\sqrt{(11,050)^{2}-4\times 1\times(-90,000)}}{2}$ (31)
-Calculating the discriminant:
-$D$ $=(11,050)^{2}+360,000$
-$=122,102,500+360,000=122,462,500$ (32)
-Finding the square root:
-$\sqrt{D}\approx 11,062.83$ (33)
-Determining the positive root:
-$\Delta N=\frac{-11,050+11,062.83}{2}\approx\frac{12.83}{2}\approx 6.415$ (34)
-The computational cost of MOTR exceeds that of FastTrackTr when the number of additional queries satisfies $\Delta N\geq 7$.
-This indicates that increasing the number of queries by at least seven will result in higher computational complexity for MOTR compared to FastTrackTr, considering all components of the decoder layer.
-In summary, the computational load of FastTrackTr is generally lower than that of MOTR, except in scenarios where there are very few objects to track; in those cases, the theoretical computation may be slightly higher for FastTrackTr.
-Otherwise, FastTrackTr consistently requires less computational effort.
-Additionally, in practical applications, MOTR tends to be significantly slower due to memory constraints.
-For a detailed analysis, please refer to Section.
-B.
-A.2 Possibilities of End-to-end
-We originally developed FastTrackTr with the goal of proposing a real-time, end-to-end multi-object tracking model.
-However, we observed that eliminating one decoder and directly applying MOTIP’s ID loss resulted in significant challenges in model convergence.
-Upon further analysis, we determined that although MOTIP’s ID loss appears similar to the appearance loss used in the previous JDT models, they fundamentally differ.
-Figure 5: The proposed architecture features a hybrid decoder capable of facilitating end-to-end rapid tracking, which currently consists of six layers.
-There are two distinct types of layers within this decoder: the standard DETR layers, which are used to output detection results, and the ID decoder layers, which are dedicated to producing tracking outcomes.
-These layers are arranged alternately, with each type comprising three layers.
-This design aims to effectively merge detection and tracking functionalities within a single streamlined framework.
-For models like JDE [37], the focus is on the similarity of object appearances, which does not incorporate specific object trajectories.
-In contrast, the ID decoder in MOTIP primarily handles trajectory information, with image data serving only as a supplementary element.
-The function and training method of this decoder might be similar to the entire PuTR model.
-Although our FastTrackTr attempts to integrate some temporal information by passing queries from one frame to the next, these queries predominantly contain image features.
-Under such conditions, supervising motion trajectories directly using MOTIP’s method becomes less effective, leading to poor or nearly impossible model convergence.
-Following this analysis, we attempted to incorporate more historical information into the decoder.
-Our preliminary design introduces an identity-specific decoder layer into the traditional DETR architecture, as illustrated in Figure.5.
-These layers enhance alignment between current frame detections and corresponding identity features through a two-step attention mechanism, and link current detections with historical trajectories by emphasizing temporal consistency.
-The ID decoder layers are interlaced with standard DETR layers, initially comprising six layers in total.
-The trajectory attention within the ID decoder layers can be formally described as follows:
-[…]
-where $emb_{t}$ and the input id decoder in MOTIP are the same.
-Its main components are the output of the previous decoder layer and the id encoding.
-$emb_{h}$ is the historical encoding.
-Finally, the $q^{\prime}_{ID}\in\mathbb{R}^{2\times N\times C}$ will be reduced in dimension and fused through a conv layer and an FFN layer in the feedforward network.
-The id loss used after the feedforward neural network and the id head is the id loss of MOTIP, which is shown below:
-$&\mathcal{L}_{id}=\frac{-\sum_{t=2}^{T+1}\sum_{m=1}^{M_{t}}\sum_{k=1}^{K+1}y_{m}^{k}\log\left(p_{m}^{k}\right)}{\sum_{t=2}^{T+1}M_{t}},\\ &y_{m}^{k}=\begin{cases}1&\mathrm{gt}_{id}\text{ of }m^{th}\text{ object is }k,\\ 0&\text{ else },\end{cases}\end{aligned}$ (36)
-where $T+1$ denotes the number of frames in a training clip, while there are $M_{t}$ ground-truth objects in the $t$-th frame.
-$y_{m}^{k}$ is an indicator function according to the identity ground truth of each object, as shown in Eq.(37) .
-In practice, we simultaneously train object detector and ID predictor in an end-to-end strategy.
-Therefore, we leverage an overall loss function $\mathcal{L}$ to supervise these two parts:
-$\mathcal{L}=\lambda_{\text{det}}\mathcal{L}_{\text{det}}+\lambda_{\text{id }}\mathcal{L}_{\text{id }}$ (37)
-Due to time constraints and the incomplete state of the model, we have conducted only one epoch of training on the DanceTrack dataset to observe whether the loss converges.
-After this single epoch, when employing the MOTIP process during inference, the HOTA score was approximately 16.7.
-This result preliminarily confirms that the model is broadly effective, with further fine-tuning required to optimize its performance.
-A.3 Improvements of the JDE association module
-We believe that one of the reasons why JDT-type models have gradually been abandoned in recent years is that the commonly used JDE [37] association module has become somewhat outdated.
-This has led to JDT models underperforming compared to some of the more recent models.
-To better demonstrate the superiority of our approach, we made some minor adjustments to the JDE association method to improve its tracking performance.
-Our improvements mainly focus on two aspects:
-  • 1)
-    In the JDE association module, we associate only high-confidence detection boxes, ignoring those with lower confidence scores.
-By incorporating ByteTrack’s [48] bi-stage matching mechanism, we temporarily retain these low-confidence detection boxes as candidate targets.
-After matching the high-confidence targets, we perform a second round of matching using the high-confidence targets, which enhances tracking performance to some extent.
-  • 2)
-    We also made some adjustments to the association strategy.
-In JDE, the first association is typically based on ID features, followed by a second association using historical trajectories.
-In our approach, similar to many other trackers, we include historical trajectories in the first round of association.
-Specifically, the cost matrix $C$ is the weighted sum of the appearance cost $A_{a}$ and the motion cost $A_{m}$, as follows:
-Figure 6: The changes in GPU memory consumption after the decoder for the MOTR, Deformable DETR, and our improved model in MOT17-03.
-Figure 7: A simple MOTR and our improved tracking process diagram. The number of queries for MOTR will change, but after our improvement, it remains unchanged.
-$C=\lambda A_{a}+(1-\lambda)A_{m}$ (38)
-where the weight factor $\lambda$ is set to 0.99 , as in [21, 43] .
-Table 9: Ablation Study of association module on the DanceTrack validation set.
-Byte Cost HOTA DetA AssA IDF1
-51.5 72.4 37.5 51.4
-$\checkmark$ 53.4 74.3 39.7 53.1
-$\checkmark$ 52.4 73.1 38.3 52.6
-$\checkmark$ $\checkmark$ 54.1 75.6 40.0 54.7
-Table 10: Impact of different of positional encodings of the historical encode on the DanceTrack validation set.
-Mechanism HOTA DetA AssA IDF1
-w///o 53.8 74.9 39.5 53.8
-MLP 53.5 74.2 38.9 53.5
-decoder 54.1 75.6 40.0 54.7
-A.4 More Implementation Details
-By default, on DanceTrack [34], we train FastTrackTr for 27 epochs on the training set, dropping the learning rate by a factor of 10 at the 14th and 18th epochs.
-On SportsMOT [7], we train our model for 27 epochs on the training set, with the learning rate dropped at the 14th and 18th epochs.
-On MOT17 [23], we train the model on this combined training set for 80 epochs, with the learning rate dropped at the 45th and 55th epochs.
-For the additional use of the CrowdHuman dataset on DanceTrack and MOT17, our usage method follows the same approach as MOTIP. Interested readers can refer to the original work for further details.
-Lastly, regarding the positional encoding of the historical encoder, the Table 10 shows the impact of different positional encodings on performance.
-In this table, ”w/o” denotes the absence of positional encoding, ”MLP” indicates that the positional encoding is generated from an MLP, and ”decoder” refers to using the positional encoding from the final layer of the decoder.
-Appendix B How to Improve MOTR inference Speed
-[…]
-This led us to speculate whether the issue might be due to the dynamic nature of the query system, which could result in variable memory usage.
-As shown in Figure 6, we compare the memory consumption of the decoder when using fixed queries versus dynamic queries.
-After further investigation, we found that this is indeed a plausible cause.
-In cuda, memory is managed through cudaMalloc and cudaFree.
-These operations incur significant computational overhead.
-To mitigate this, PyTorch attempts to reuse memory blocks previously allocated via cudaMalloc.
-When a suitable cached block exists in PyTorch’s allocator, it directly serves the request without invoking cudaMalloc, thereby minimizing latency during initial allocation phases.
-However, MOTR’s dynamic data patterns require intermediate tensors of varying sizes throughout runtime.
-This variability prevents the allocator from leveraging cached blocks of appropriate dimensions, forcing frequent calls to cudaFree to release outdated allocations and subsequent cudaMalloc requests for new memory spaces.
-Consequently, the allocator repeatedly rebuilds its caching infrastructure through bulk CUDA API operations – an expensive process that introduces substantial latency spikes.
-This thrashing behavior substantially degrades computational efficiency and hinders full GPU utilization
-We experimented with PyTorch’s memory management code, and observed that when the tensor size remains relatively constant, there are significantly fewer memory allocation and release calls.
-The specific code we tested can be easily found in GitHub, when you search for ’torch mem’(we would like to note that the owner of that link is not associated with the authors of this paper.
-So how can we address this issue?
-Modifying PyTorch’s underlying logic directly is quite complex and would likely require intervention from PyTorch’s developers (we are preparing to submit an issue on GitHub or send an email to the maintainers).
-An alternative approach could be to fix the number of tracking queries in MOTR and use masks to differentiate between empty tracking queries and those that are actively used.
-This may resolve the issue.
-As shown in Figure 7, we trained an updated version of MOTR for one epoch on the DanceTrack dataset and tested the speed and GPU utilization.
-Due to limitations in funding, equipment, and time, we did not fully train the model to evaluate its performance.
-However, preliminary results show promise: after one epoch, the HOTA score on the validation set reached 18.5 and fps has increased from 16.5 to 25 with fp32 precision.
-Note that we have not reduced any network structure at this time.
-Interested readers are welcome to experiment with this approach themselves.
-It is relatively straightforward to implement, though we advise against using GPUs with less than 24GB of memory, as this may lead to out-of-memory errors.
-Of course, this is just one possible explanation.
-Other factors may also contribute to the underutilization of GPU resources in MOTR-like models.
-If any readers have alternative insights, we welcome further discussion and collaboration.
-Appendix C Visualization of FastTrackTr Results
-In this section, we will provide visualization results of our model on different datasets.
-Figure 8: Our model’s visualization results on Dancetrack, where the same color represents the same object.
-Figure 9: Our model’s visualization results on MOT17, where the same color represents the same object.
-Figure 10: Our model’s visualization results on SportsMOT, where the same color represents the same object.